### v1

# Propensity Score Alignment of Unpaired Multimodal Data

Abstract
Multimodal representation learning techniques typically rely on paired samples to learn common representations, but paired samples are challenging to collect in fields such as biology where measurement devices often destroy the samples. This paper presents an approach to address the challenge of aligning unpaired samples across disparate modalities in multimodal representation learning. We draw an analogy between potential outcomes in causal inference and potential views in multimodal observations, which allows us to use Rubin’s framework to estimate a common space in which to match samples. Our approach assumes we collect samples that are experimentally perturbed by treatments, and uses this to estimate a propensity score from each modality, which encapsulates all shared information between a latent state and treatment and can be used to define a distance between samples. We experiment with two alignment techniques that leverage this distance—shared nearest neighbours (SNN) and optimal transport (OT) matching—and find that OT matching results in significant improvements over state-of-the-art alignment approaches in both a synthetic multi-modal setting and in real-world data from NeurIPS Multimodal Single-Cell Integration Challenge.
1 Introduction
Large-scale multimodal representation learning techniques such as CLIP (Radford et al., 2021) have lead to remarkable improvements in zero-shot classification performance and have enabled the recent success in conditional generative models. However, the effectiveness of multimodal methods hinges on the availability of paired samples—such as images and their associated captions—across data modalities. This reliance on paired samples is most obvious in the InfoNCE loss (Gutmann & Hyvärinen, 2010; van den Oord et al., 2018) used in CLIP (Radford et al., 2021) which explicitly learns representations to maximize the true matching between images and their captions.
While paired image captioning data is abundant on the internet, paired multimodal data is often challenging to collect in scientific experiments. Take, for instance, experiments in biology, where unpaired data are the norm for technical reasons: RNA sequencing, protein expression assays, and the collection of microscopy images for cell painting experimental assays are all destructive processes. Because of this, we cannot collect multiple different measurements from the same cell, and can only explicitly group cells by their experimental condition.
If we could accurately match unpaired samples across modalities, we could use the aligned samples as proxies for paired samples and apply existing multimodal representation learning techniques. However, matching across modalities is challenging for two reasons. Firstly, observations from various modalities often exist in entirely different spaces. For instance, microscopy images of cells are in pixel space, while gene expression assays provide data in the form of mRNA abundance counts. Secondly, even if measurements were taken in the same space, determining an appropriate distance metric for matching remains a non-trivial task. For example, when matching images of distinct cells in different microscopy images, we would ideally want a metric that is more sensitive to features describing the underlying state of the cell than features related to biologically-irrelevant cell appearance, such as the orientation of the cell.
To understand the matching problem more abstractly, we can think of the multiple modalities as potential “views”, $x^{(1)}(z)\in\mathcal{X}^{(1)},x^{(2)}(z)\in\mathcal{X}^{(2)}$ of the same underlying latent state, $z\in\mathcal{Z}$, where we only get to observe one view for any individual unit $i$ (e.g. an individual cell in our motivating example). Observing, or accurately inferring $z$ would address the first issue by providing a common space $\mathcal{Z}$ in we could match samples. However, inferring the latent variable is hopelessly underspecified without making strong assumptions on the data generating processes of both modalities (Xi & Bloem-Reddy, 2023). Furthermore, for complex systems such as biological samples, $z$ may still be extremely high-dimensional, and as a result, the second challenge remains: what is the “right” metric is for measuring similarity between samples in this space?
We address both challenges in this paper by appealing to classical ideas from Rubin (1974), in the case where we additionally observe a label for each unit $t\in\mathcal{T}$, e.g., indexing an experiment that it belongs to. By making the assumption that $t$ perturbs the observations via their shared latent state, we identify an observable link between modalities with the same underlying $z$. Our key observation is that the propensity score, defined as $p(t|z)$, satisfies three properties: 1) It provides a common space for matching, 2) it maximally coarsens $z$, retaining only relevant information and 3) under certain assumptions, it is estimable from observations of individual modalities alone (Proposition 3.1).
Figure 1: Our problem setting and proposed solution. We assume that two observed modalities (e.g., Microscopy images and RNA expression data) are generated in pairs according to some common latent variable, $z_{i}$, but we only observe the respective modalities separately. We further assume the latents are experimentally perturbed by a treatment, $t$, and we get observations from both modalities from a control state and at least one interventional state. Under our assumptions, training separate classifiers that predict the intervention label reveals the shared information $p(t|z_{i})$, which allows us to re-pair the observed disconnected modalities.
In practice, the propensity score is straightforward to estimate: one simply trains two classifiers—one for each modality—to predict which treatment was applied to the respective modalities, and then match across modalities based on the similarity between predicted treatments within each treatment group. The proposed methodology is generic across modalities—it can be applied to match observations between any modality for which a classifier can be efficiently trained. However, we cannot naively use bipartite matching, as the same sample unit does not appear in both modalities. To address this, we use soft matching techniques to estimate the missing modality for each sample unit by combining multiple observations. We experiment with two different matching approaches from the recent literature: shared nearest neighbours (SNN) matching (Lance et al., 2022; Cao & Gao, 2022) and optimal transport (OT) matching Villani (2009); Schiebinger et al. (2019), and find that OT matching with distances defined on the propensity score lead to significant improvement on real world multi-modal matching tasks from the NeurIPS Multimodal Single-Cell Integration Challenge Lance et al. (2022). We found similarly large improvements on synthetic image-based benchmarks.
Finally, we show how our matching procedure can be used for cross-modality prediction tasks. We estimate a function that translates between modalities by integrating over the coupling matrix output by our alignment procedure. Our approach uses a two-sample stochastic gradient estimator to get an unbiased estimate of the gradient of the resulting loss. In our experiments, this approach gave very strong performance, even outperforming methods that had access to the ground truth matching.
2 Related work
Unpaired Data   Dealing with unpaired data is important in translation tasks, either for images (Liu et al., 2017; Zhu et al., 2017; Almahairi et al., 2018), or more recently for biological modalities (Amodio & Krishnaswamy, 2018; Yang et al., 2021). In particular, Yang et al. (2021) formalize this setting as being generated by a shared underlying latent variable, which has also been identified as a useful modelling assumption for identifiability analyses in multi-view or multi-modal representation learning (Gresele et al., 2020; Sturma et al., 2023). Our work also exploits this modelling assumption for theoretical justification, but, unlike previous works, our method does not depend on performing inference over the latent space, thus avoiding the identifiability issue. Specifically, since we do not require modality translation, we do not require an encoder-decoder structure (Yang et al., 2021), which embeds modality-specific reconstruction information in the latent space that can be detrimental for the simpler matching task.
Optimal Transport Matching   Optimal transport is commonly used to solve a similar problem in single-cell biology known as cell trajectory inference. In this case, $X^{(1)}$ to $X^{(2)}$ are random variables measured at different time points in a shared (metric) space, and OT matching can be achieved by minimizing the metric (Schiebinger et al., 2019; Tong et al., 2020). Recent work (Demetci et al., 2022) tackles our multi-modal alignment problem by using the Gromov-Wasserstein distance, which leverages local metric structure within pairs of points from each modality (Demetci et al., 2022). In addition to this “pure” OT approach, Gossi et al. (2023) use OT on contrastive learning representations, though this approach requires matched pairs for training, while Cao et al. (2022) use OT in the latent space of a multi-modal VAE.
Perturbations and Heterogeneity   Many methods treat observation-level heterogeneity as another dimension to globally integrate, even when cluster labels are observed (Butler et al., 2018; Korsunsky et al., 2019; Foster et al., 2022). This is sensible when clusters correspond to noise rather than the signal of interest. However, it is well known in causal representation learning that heterogeneity—particularly heterogeneity arising from different perturbations—has theoretical benefits (Khemakhem et al., 2020; Squires et al., 2023; Ahuja et al., 2023; Buchholz et al., 2023; von Kügelgen et al., 2023). There, the benefits (weakly) increase with the number of perturbations, which is also true of our setting (Proposition 3.2). In the context of aligning unpaired data, only Yang et al. (2021) leverage this heterogeneity in their method. Specifically, they require VAE representations to classify experimental labels in addition to reconstructing modalities, while our method only requires a representation to classify labels. Yang et al. (2021) treat the classifier objective as a regularizer, but our theory suggests that it is responsible for their matching performance. In our experiments, we found that requiring reconstruction led to worse matching performance with equivalent architectures.
3 Propensity Score Multi-modal Matching
We consider multimodal settings where there exist two potential views, $X^{(e,t)}\in\mathcal{X}^{(e)}$ from two different modalities indexed by $e\in\{1,2\}$ and experiment $t$ that perturbs a shared latent state of these observations. The observations will typically be in very different spaces: for example, $\mathcal{X}^{(1,t)}$ may be the space of images of cells under a microscope, and $\mathcal{X}^{(2,t)}$ may be the space of gene expression data. This process defines a jointly distributed random variable $(X^{(1,t)},X^{(2,t)},e,t)$, from which we observe only a single modality, its index, and treatment assignment,$\{x_{i}^{(e_{i},t_{i})},e_{i},t_{i}\}_{i=1}^{n}$ (we will always denote random variables by upper-case letters, and samples by their corresponding lower-case letter). Our aim is to match or estimate the samples from the missing modality that would correspond to the realization of the missing random variable. Importantly, we match within treatment groups, and the treatment variable is only used to learn a common space in which to match.
We assume each modality arises from a common latent random variable $Z$ as follows,
$t\sim P_{T},\quad Z^{(t)}\mid t\sim P_{Z}^{(t)},\quad U^{(e)}\sim P_{U}^{(e)},\quad U^{(e)}{\perp\!\!\!\perp}Z,$
[…]
Under this model, if $Z$ were observable, an optimal matching can be constructed by simply matching the modalities with the most similar $Z$. However, $Z$ is latent, and inference on the model described by Section 3 is arguably more difficult than the matching problem itself due to theoretical difficulties such as identifiability and disentangling $Z$ from the modality-specific noise terms $U^{(e)}$.
Instead, we see that the interventions $t$ provide an observable link between the modalities, thereby revealing information about $Z^{(t)}$. Specifically, we use the propensity score with respect to $t$, which we define as
$\pi(z):=P(t|Z=z)\in[0,1]^{T+1},$ (2)
as a proxy for the latent $Z$. Now, although we cannot compute this directly as $Z$ is latent, we make the observation that if $f^{(e)}$ is injective for $e=1,2$, then we can compute the compute the propensity score from each of the observed modalities, since it will be that $\pi(Z^{(t)})=\pi(X^{(e,t)})$, for both $e=1$ and $e=2$. Not only does the propensity score reveal shared information, classical causal inference (Rubin, 1974) states that it captures all information shared between the latent and treatment, and does so minimally, in terms of having minimum dimension and entropy. We collect these observations into the following proposition.
Proposition 3.1.
In the model described by Section 3, further assume that $f^{(e)}$ are injective for $e=1,2$. Then, the propensity score in either modality is equal to the propensity score given by $Z^{(t)}$, i.e., $\pi(X^{(1,t)})=\pi(X^{(2,t)})=\pi(Z^{(t)})$ as random variables. This implies
$I(t,Z^{(t)}\mid\pi(Z^{(t)}))=I(t,Z^{(t)}\mid\pi(X^{(t)}))=0,$ (3)
for each $t$, where $I$ is the mutual information. Furthermore, any other function $b(Z^{(t)})$ satisfying $I(t,Z^{(t)}\mid b(Z^{(t)}))=0$ is such that $\pi(Z^{(t)})=f(b(Z^{(t)}))$.
The proof can be found in the Appendix. The above shows that computing the propensity score on either modality is equivalent to computing it on the unobserved shared latent, which captures all the shared information observable in $t$. The final statement implies that it is of minimal dimension and entropy, and thus it discards the modality-specific information that may be counterproductive to matching.
Number of Perturbations   Note that point-wise equality of the propensity score $\pi(z_{1})=\pi(z_{2})$ does not necessarily imply equality of the latents $z_{1}=z_{2}$, due to potential non-injectivity. In fact, suppose that $Z^{(t)}$ is supported in $\mathbb{R}^{d}$. The propensity score $\pi(z)$ maps into the $T$-dimensional simplex, where $T$ is the number of perturbations, which can be parameterized by coordinates in $\mathbb{R}^{T}$ (essentially, the logits of the classifier). Under mild conditions, the propensity score necessarily compresses information about $Z^{(t)}$ if the latent dimension exceeds the number of perturbations, echoing impossibility results from the causal representation learning literature (Squires et al., 2023).
Proposition 3.2.
Suppose that $P_{Z}^{(t)}$ has a smooth density $p(z|t)$ for each $t=0,\dots,T$. Then, if $T<d$, the propensity score $\pi$, restricted to its strictly positive part, is non-injective.
The proof can be found in the Appendix. Note the above only states an impossibility result when $T<d$. When $T\geq d$, it can be seen from the proof of Proposition 3.2 that the injectivity of the propensity score depends on the injectivity of the following expression in $z$:
$f(z)=\begin{bmatrix}\log(p(z|t=1))-\log(p(z|t=0))\\ \vdots\\ \log(p(z|t=T))-\log(p(z|t=0))\end{bmatrix},$ (4)
which then depends on the latent process itself. If the above is non-injective, this is a fundamental indeterminacy that cannot be resolved without making strong assumptions on point-wise latent variable recovery, as we have already seen that the propensity score contains the maximal shared information, as shown in Proposition 3.1. Nonetheless, note that $f$ in Equation 4 is injective if any of the subset of its entries are, clearly outlining the benefits of collecting data from a larger number of perturbations for matching.
Key Assumptions   Our results thus far rely on two key modelling assumptions,
  • (A1): $t$ has a non-trivial effect on $Z$, but does not affect $u^{(e)}$, implying that interventions are able to target the common underlying process without changing modality-specific properties.
  • (A2): Injectivity of $f^{(e)}$ in the propensity score derivation, $f^{(e)}(z,u)=f^{(e)}(z^{\prime},u^{\prime})\Rightarrow(z,u)=(z^{\prime},u^{\prime})$, which implies each modality contains complete information about the underlying process.
(A1) is justifiable insofar as modalities represent different measurements of an isolated system, such as in biological studies where $Z$ might correspond to an underlying cell state and modalities refer to different single-cell measurements. However, in practice, even different measurement devices may be more or less sensitive to the biological variation implied by $t$. We weaken this assumption in a simplified case in Section 3.
(A2) requires that no information about the latent state is lost in the observation process. Note that the injectivity is in the sense of $f$ as a function of both $u$ and $z$, which allows observations that have a shared $z$ but differ by their value in $u$, and the function remains injective. For example, rotated images with the exact same content can have a shared $z$, but remain injective due to the rotation being captured in $u$.
Relaxing Assumption 1
Consider the propensity score
$\pi(x^{(e,t)})=P(t|X^{(e,t)}=x^{(e,t)})$ (5)
where we do not necessarily require $U^{(e)}{\perp\!\!\!\perp}t\mid Z^{(t)}$, and thus we obtain
$\pi(x^{(1,t)})=P(t|Z^{(t)}=z^{(t)},U^{(1)}=u^{(1)})\neq$ (6)
$P(t|Z^{(t)}=z^{(t)},U^{(2)}=u^{(2)})=\pi(x^{(2,t)}),$
see the proof of Proposition 3.1 for details.
Suppose that the two observed modalities are indeed generated by a shared $\{z_{i}\}_{i=1}^{n}$, but where the indices of modality $2$ are potentially permuted, and with values differing by modality specific information:
$\{x^{(1,t)}_{i}=f^{(1)}(z_{i},u_{i}^{(1)})\}_{i=1}^{n},\{x^{(2,t)}_{j}=f^{(2)}(z_{2},u_{j}^{(2)})\}_{j=1}^{n},$ (7)
where $j=\pi(i)$ denotes a permutation of the sample indices. Under (A1), we would be able to find some $j$ such that $\pi(x^{(1,t)}_{i})=\pi(x^{(2,t)}_{j})$ for each $i$.
Matching via OT can allow us to relax (A1) in a very particular way. Consider the simple case where $t\in\{0,1\}$, so that $\pi$ can be written in a single dimension, e.g., $P(t=1|X^{(e,t)}=x^{(e,t)})\in[0,1]$. In this case, exact OT is equivalent to sorting $\pi(x^{(1,t)}_{i})$ and $\pi(x^{(2,t)}_{j})$, and matching the sorted versions 1-to-1. Under (A1), the sorted versions will be exactly equal. A relaxed version of (A1) that would still result in the correct ground truth matching is to assume that $t$ affects $U^{(1)}$ and $U^{(2)}$ differently, but that the difference is order preserving, or monotone. Denote $(\pi(x^{(1,t)}_{i}),\pi(x^{(2,t)}_{i}))=(\pi_{i}^{(1)},\pi_{i}^{(2)})$ as the true pairing, noting that we use the same index $i$. We require the following:
$(\pi_{i_{1}}^{(1)}-\pi_{i_{2}}^{(1)})(\pi_{i_{1}}^{(2)}-\pi_{i_{2}}^{(2)})\geq 0,\quad\forall i_{1},i_{2}=1,\dots,n.$ (8)
This says that, even if $\pi_{i}^{(1)}\neq\pi_{i}^{(2)}$, that their relative orderings will still coincide. Then, exact OT will still recover the ground truth matching. See Figure 2 for a visual example of this type of monotonicity. For example, suppose that $t$ is a chemical perturbation of a cell, and thus $\pi_{i}^{(1)}$, $\pi_{j}^{(2)}$ can be seen as a measure of biological response to the perturbation, e.g., in a treated population, $\pi_{i_{1}}>\pi_{i_{2}}$ indicates samples $i_{1}$ had a stronger response than sample ${i_{2}}$, as perceived by the first modality indexed by $i$. Then, this monotonocity states that we should see the same $\pi_{j_{1}}>\pi_{j_{2}}$ in the other modality as well, if the samples $i_{1}$ and $i_{2}$ truly corresponded to $j_{1}$ and $j_{2}$.
Figure 2: OT matching allows for $t$ to have different effects on the modality specific information, here $u_{i}^{(1)}$ and $u_{i}^{(2)}$, as long as they can be written as transformations that preserve the relative order within modalities. Exact OT in 1-d always matches according to the relative ordering, and thus exhibits this type of “no crossing” behaviour shown in the figure on the left. The figure on the right shows a case where we would fail to correctly match across modalities because of the crossing shown in orange.
When $t$ is not a binary treatment and the propensity scores are multidimensional, a condition known as cyclic monotonicity is sufficient for recovering the true matching, see Appendix A for details. However, unlike the 1-d case, we are not aware of more intuitive conditions under which we should expect cyclic monotonicity to hold.
4 Methodology
For the remainder, we will drop $(e,t)$ from our notation, and use $(x_{i},t_{i})$ to denote observations from modality 1, and $(x_{j},t_{j})$ to denote observations from modality 2 whenever it does not cause confusion to do so. Given a multimodal dataset with observations $\{(x_{i},t_{i})\}_{i=1}^{n_{1}}$ and $\{(x_{j},t_{j})\}_{j=1}^{n_{2}}$, we wish to compute a matching matrix (or coupling) between the two modalities. We define a $n_{1}\times n_{2}$ matching matrix $M$ where $M_{ij}$ represents the likelihood of $x_{i}$ being matched to $x_{j}$. We always perform matching only within observations with the same value of $t$, so that in practice we obtain a matrix $M_{t}$ for each $t$.
Our method approximates the propensity scores by training separate classifiers that predicts $t$ given $x$ for each modality. We denote the estimated propensity score by $\pi_{i}$ and $\pi_{j}$ respectively, where
$\pi_{i}\approx\pi(x_{i})=P(T=t\mid X_{i}^{(e,t)}=x_{i}).$ (9)
This results in the transformed datasets $\{\pi_{i}\}_{i=1}^{n_{1}}$ and $\{\pi_{j}\}_{j=1}^{n_{2}}$, where $\pi_{i}$, $\pi_{j}$ are in the $T$ dimensional simplex. We use this correspondence to compute a cross-modality distance function:
$d(x_{i},x_{j}):=d^{\prime}(\pi_{i},\pi_{j}),$ (10)
where in practice, we typically compute the Euclidean distance in $\mathbb{R}^{T}$ of the logit-transformed classification scores. Given this distance function, we rely on existing matching techniques to constructing a matching matrix. In our experiments, we found OT matching gave the best performance, but we also evaluated Shared Nearest Neighbour matching; details of the latter can be found in Appendix B.
4.1 Optimal Transport Matching
The propensity score distance allows us to easily compute a cost function associated with transporting mass between modalities, $c(x_{i},x_{j})=d^{\prime}(\pi_{i}\pi_{j})$. Let $p_{1},p_{2}$ denote the uniform distribution over $\{\pi_{i}\}_{i=1}^{n_{1}}$ and $\{\pi_{j}\}_{j=1}^{n_{2}}$ respectively. Discrete OT aims to solve the problem of optimally redistributing mass from $p_{1}$ to $p_{2}$ in terms of incurring the lowest cost. Let $C_{ij}=c(x_{i},x_{j})$ denote the $n_{1}\times n_{2}$ cost matrix. The Kantorovich formulation of optimal transport aims to solve the following constrained optimization problem:
$\min_{M}\sum_{i}^{n_{1}}\sum_{j}^{n_{2}}C_{ij}M_{ij},\quad M_{ij}\geq 0,\quad M\mathbf{1}=p_{1},\quad M^{\top}\mathbf{1}=p_{2}.$
This is a linear program, and for $n_{1}=n_{2}$, it can be shown that the optimal solution is a bipartite matching between $\{\pi_{i}\}_{i=1}^{n_{1}}$ and $\{\pi_{j}\}_{j=1}^{n_{2}}$. We refer to this as exact OT; in practice we add an entropic regularization term, resulting in a soft matching, that ensures smoothness and uniqueness, and can be solved efficiently using Sinkhorn’s algorithm. Entropic OT takes the following form:
$\min_{M}\sum_{i}^{n_{1}}\sum_{j}^{n_{2}}C_{ij}M_{ij}-\lambda H(M),\quad M_{ij}\geq 0,\quad M\mathbf{1}=p_{1},\quad M^{\top}\mathbf{1}=p_{2},$
where $H(M)=-\sum_{i,j}M_{ij}\log(M_{ij})$, the entropy of the joint distribution implied by $M$. This approach regularizes towards a higher entropy solution, which has been shown to have statistical benefits (Genevay et al., 2018), but nonetheless for small enough $\lambda$ serves as a computationally appealing approximation to exact OT.
5 Cross-modality prediction
Given a matching matrix $M$, we can interpret this as the probability that each sample, $i$, from modality (1) is matched to sample $j$ in modality (2); that is $M_{i,j}=P(x^{(2)}_{j}|x^{(1)}_{i})$. As a downstream task, and as an evaluation metric for $M$, we can use this matching to estimate a cross-modal prediction model, $f_{\theta}$, that maps from modality (2) to (1) by minimizing the following loss,
$\mathcal{L}(\theta):=\sum_{i}(x_{i}^{(1)}-M_{i}f_{\theta}(x_{j}^{(2)}))^{2}.$ (11)
However, this requires evaluating $f_{\theta}$ for all $n_{2}$ examples from modality (2) for each of the $n_{1}$ examples in modality (1). Of course, we can avoid this cost with stochastic gradient descent by sampling from modality $(2)$ via $M_{i\cdot}$ for each training example $(1)$, but to get an unbiased estimate of $\nabla_{\theta}\mathcal{L}$, we need two samples from modality (2) for each sample from modality (1),
$\nabla\mathcal{L}(\theta)\approx$ $-2\left(x_{i}^{(1)}-f_{\theta}(\dot{x}_{j}^{(2)})\right)\nabla_{\theta}f_{\theta}(\ddot{x}_{j}^{(2)})\quad$ $\dot{x}_{j}^{(2)},\ddot{x}_{j}^{(2)}\sim P(x^{(2)}_{j}|x^{(1)}_{i}).$ (12)
By taking two samples as in equation (12), we get an unbiased estimator of $\nabla\mathcal{L}(\theta)$, whereas a single sample would have resulted in optimizing an upper-bound on equation (11); for details, see Hartford et al. (2017) where a similar issue arises in the gradient of their causal effect estimator.
6 Experiments
We evaluate our proposed methodology on two datasets. The first is a synthetic interventional image dataset generated satisfying the assumptions of Section 3. The second is a real-world single-cell CITE-seq data obtained from the NeurIPS 2021 Multimodal single-cell data integration competition (Lance et al., 2022), which provides a ground truth matching by allowing for a small number of cell surface proteins to be measured simultaneously to RNA sequencing. Note the CITE-seq dataset is not interventional—we use the cell type as the classification target $t$ instead. All details are made available in the Appendix. In both cases the ground truth matching is known to make evaluation possible, but hidden during training.
Baselines   In terms of the problem setting, the most closely related method is Gromov-Wasserstein optimal transport (SCOT) (Demetci et al., 2022), which explicitly computes a cross-modality cost for optimal transport matching by leveraging the local geometry within each modality. For the CITE-seq data, graph-linked VAE (scGLUE), which has access to specific biological metadata that connects the biological modalities, can also be used for matching. [^2] A variant of scGLUE was the top entrant in the NeurIPS 2021 Multi-modal single-cell data integration competition, from which we source our data. However, ground truth matchings were made available in the training set in that case, which fundamentally changes the nature and difficulty of the problem. However, neither of the preceding approaches use the experimental label information, which is a key component of our method. In fact, the only method to our knowledge that leverages such information is the VAE approach developed in Yang et al. (2021), which also minimizes a classification loss from the latent space, which we will consider our main baseline. For all methods besides SCOT, we use both SNN and OT to perform the final matching in the latent space. Note that when combined with OT, our VAE baseline also resembles Cao et al. (2022).
Figure 3: VAE and classifier validation metrics on the CITE-seq dataset. Notice that validation cross-entropy inversely tracks the ground truth matching metrics, and thus can be used as a proxy in practical settings where the ground truth is unknown. The same pattern does not hold for the VAE (Yang et al., 2021), which we suspect is because reconstruction is largely irrelevant for matching.
Model Details   For our method and the VAE, we use the general architecture of a linear classification head on top of a suitable encoder. The architecture of this encoder is always the same between the VAE and our method. For the image dataset, we use the convolutional neural network from Yang et al. (2021) as the encoder, and for the CITE-seq dataset, we use fully-connected MLPs for both RNA (top 200 PC’s) and protein data. Optimization was performed using Adam with a one-cycle learning rate scheduler. We use the standard cross-entropy loss to train our classifiers for propensity score estimation. For other methods, we use existing implementations with suggested default settings.
Matching Details   Both SNN and OT use the Euclidean distance function to determine neighbours and compute the cost matrix, respectively. SCOT uses the correlation distance by default, and we found that this resulted in better performance than Euclidean distance. We use only a single neighbour for SNN matching, which interestingly resulted in the best performance. Both SCOT and OT solve the entropic regularized OT, for these we use a regularization parameter of $0.05$.
Method MSE Trace
(Med (Q1, Q3)) (Med (Q1, Q3)) ×$\times 10^{-3}$
SCOT 0.0354 0.5964
VAE+SNN 0.0622 3.116
(0.0571, 0.0676) (2.818, 3.213)
VAE+OT 0.0324 7.733
(0.0316, 0.0350) (7.473, 7.794)
PS+SNN 0.0552 7.924
(0.0530, 0.0558) (7.569, 9.504)
PS+OT 0.0316 18.329
(0.0300, 0.0330) (17.068, 18.987)
Rand 0.0709 N/A
(0.0707, 0.0714)
Table 1: Results on synthetic image data. Metrics are computed on a held out test set over 12 groups corresponding to interventions on the latent position, with approximately 1700 observations per group. A trace of $10$ for example corresponds to a total weight of $1710\times 10^{-3}\times 1700=$ 17 on the true matching, out of 1700.
Method FOSCTTM Trace
(Median (Q1, Q3)) (Median (Q1, Q3))
SCOT 0.4596 0.0200
GLUE+SNN 0.4412 0.0362
GLUE+OT 0.5309 0.0323
VAE+SNN 0.3816 0.0612
(0.3760, 0.3822) (0.0588, 0.0634)
VAE+OT 0.3953 0.0814
(0.3912, 0.4045) (0.0777, 0.8895)
PS+SNN 0.3126 0.0941
(0.3121, 0.3160) (0.0880, 0.0989)
PS+OT 0.3049 0.1163
(0.3008, 0.3078) (0.1093, 0.1250)
Table 2: Results on CITE-seq data. Metrics are computed on a held out test set averaged over a total of 45 cell types (groups) with a varying amount of observations per group. As such, the average trace can be difficult to interpret (for example, the maximum trace is near $1$, corresponding to a near-perfect matching within a group with only $7$ cells in the test set), but we remark that OT matching on propensity scores outperforms the other methods within each group as well as on average.
Evaluation Metrics   We report three evaluation metrics. The trace metric computes the average mass that the matching matrix places on the true matches (higher is better), and FOSCTTM reports the Fraction Of Samples Closer Than the True Match (FOSCTTM) ((Demetci et al., 2022), (Liu et al., 2019)) (lower is better, 0.5 corresponds to random guessing). For synthetic images where we know the ground true latent, we report the MSE to the true latent after matching. Full details of these metrics are given in Section D.1.
Modality Prediction   We also consider one metric that examines whether matched samples are useful for downstream tasks. For this, we chose the cross-modality prediction task of predicting protein levels from RNA expression in the CITE-seq dataset. To do this, we train a supervised learning model (in our case, a 2-layer MLP with either MSE loss or the unbiased gradients loss in Section 5) using samples generated by different matching matrices $M$. As baselines, we also train on pairs from the ground truth matching ($M_{ii}=1$) and random sampling ($M_{ij}=1/n$). In view of the performance observed in Table 1 and Table 2, we only considered $M$ resulting from matching with OT using the propensity score and VAE embeddings.
Method $R^{2}$ (MSE) $R^{2}$ (Unbiased)333 This refers to training the predictor with the unbiased MSE (Equation 11). The evaluation is still using $R^{2}$ as defined by the standard MSE.
(Med (Q1, Q3)) (Med (Q1, Q3))
Rand 0.1383 0.1727
(0.1372, 0.1402) (0.1701, 0.1731)
VAE+OT 0.1493 0.1142
(0.1179, 0.1724) (0.0786, 0.1594)
PS+OT 0.2174 0.2331
(0.2062, 0.2228) (0.2069, 0.2504)
True Pairs 0.2243 N/A [^4] The standard and unbiased MSE are equivalent in this case.
(0.2234, 0.2257)
Table 3: Modality prediction results on CITE-seq data. Entries represent $R^{2}$ computed on a held-out test set consisting of ground-truth pairs, which is computed as $1-\text{MSE}(\hat{x},x)/\text{MSE}(\bar{x},x)$, where $\hat{x}$ are the predicted protein levels, and $\bar{x}$ is the mean value of the validation set. Note random sampling, especially with the unbiased gradient loss, achieves a positive $R^{2}$ against $\bar{x}$ because it samples randomly within $t$, which results in learning group-level statistics.
6.1 Validation Cross-Entropy Monitor
When the assumptions are satisfied, the success of our propensity score method hinges on how well we can learn the conditional distributions $p(t\mid X_{i}^{(e,t)})$. Thus, the validation cross entropy can be used as a tractable metric that can serve as a model selection monitor for the matching performance, which would be unknown in practice. The effectiveness of such a monitor is validated in our experiments, even on real-world CITE-seq data, where we see that a lower validation loss typically corresponds to higher matching performance which is not the case for the VAE (Figure 3). We suspect that this is precisely due to the additional requirement that the VAE minimize a reconstruction loss, in addition to the classification cross-entropy.
6.2 Results
We checkpoint the classifier and VAEs at the lowest validation loss and report their metrics on a held out test set over 10 random seeds. Note that this does not necessarily select the embedding model that exhibits optimal matching, but instead the best model to select without validating against ground truth matching metrics, see Figure 3. We further used 10 random seeds for training the MLP used for modality prediction. scGLUE is trained according to the public implementation, which uses learning rate reduction and early stopping strategies. SCOT is a non-iterative approach and thus we directly report its results.
We found that OT matching on propensity scores consistently outperforms other methods on all metrics, typically followed by SNN matching on propensity scores, or OT matching on VAE embeddings. Furthermore, both the VAE and propensity score, which leverage experimental label information, tend to perform better. This suggests that using the propensity score as embeddings for matching, and using OT to perform the final matching both independently improve matching performance. Curiously, SCOT performs well on the MSE metric for image data, but only places slightly above random in the trace metric. This indicates that it matches scenes with similar latent coordinates, without placing any significant weight on the exact ground truth. This suggests that exact matches may not be necessary for a method to be useful for downstream tasks.
Surprisingly, using the propensity score embeddings with OT matching appears to improve generalization in the modality prediction task (Table 3), when using the unbiased MSE over a model trained on ground truth pairings (indeed, we observed that the ground truth model had a lower training loss, but higher test loss). This reveals an unexpected benefit of (soft) matching: we can sample from the conditional to minimize the loss Equation 11, which, with a suitable $M$, results in improved generalization compared to the naive MSE. This hinges strongly on the quality of $M$—the $M$ resulting from the VAE embeddings results in worse generalization, and in practice the quality of the matching, at least in absolute terms, remains hidden.
7 Conclusion
This work presents a simple algorithm for aligning unpaired data from different modalities. The method is both very general—only requiring a classifier to be trained on each modality—and highly effective for matching, which we show both theoretically and empirically. As a downstream task, we demonstrate the effectiveness of the resulting matchings for cross modality prediction, which leads to better generalization than the ground truth matching on the dataset we evaluated. We suspect that this improved generalization is the result of implicitly enforcing invariance to modality specific information, but more work is needed to evaluate the conditions under which this improved generalization occurs.
Appendix A Cyclic monotonicity
We can see the monotonicity requirement Equation 8 as the monotonicity of the function with graph $(\pi_{i}^{(1)},\pi_{i}^{(2)})\in[0,1]^{2}$. In higher dimensions, we require that the “graph” satisfies the following cyclic monotonicity property (Villani, 2009):
Definition A.1.
The collection $\{(\pi_{i}^{(1)},\pi_{i}^{(2)})\}_{i=1}^{n}$ is said to be $c$-cyclically monotone for some cost function $c$, if for any $n=1,\dots,N$, and any subset of pairs $(\pi_{1}^{(1)},\pi_{1}^{(2)}),\dots,(\pi_{n}^{(1)},\pi_{n}^{(2)})$, we have
$\sum_{n=1}^{N}c(\pi_{n}^{(1)},\pi_{n}^{(2)})\leq\sum_{n=1}^{N}c(\pi_{n}^{(1)},\pi_{n+1}^{(2)}).$ (13)
Importantly, we define $\pi_{n+1}=\pi_{1}$, so that the sequence represents a cycle.
Note in our setting, the OT cost function is the Euclidean distance, $c(x,y)=\|x-y\|_{2}$. It is known that the OT solution must satisfy cyclic monotonicity. Thus, if the true pairing is uniquely cyclically monotone, we can recover it with OT.
Appendix B Shared Nearest Neighbours Matching
Using the propensity score distance, we can compute nearest neighbours both within and between the two modalities. We follow Cao & Gao (2022) and compute the normalized shared nearest neighbours (SNN) between each pair of observations as the entry of the matching matrix. For each pair of observations $(\pi_{i}^{(1)},\pi_{j}^{(2)})$, we define four sets:
  • $\texttt{11}_{ij}$: the k nearest neighbours of $\pi_{i}^{(1)}$ amongst $\{\pi_{i}^{(1)}\}_{i=1}^{n_{1}}$. $\pi_{i}^{(1)}$ is considered a neighbour of itself.
  • $\texttt{12}_{ij}$: the k nearest neighbours of $\pi_{j}^{(2)}$ amongst $\{\pi_{i}^{(1)}\}_{i=1}^{n_{1}}$.
  • $\texttt{21}_{ij}$: the k nearest neighbours of $\pi_{i}^{(1)}$ amongst $\{\pi_{j}^{(2)}\}_{j=1}^{n_{2}}$.
  • $\texttt{22}_{ij}$: the k nearest neighbours of $\pi_{j}^{(2)}$ amongst $\{\pi_{j}^{(2)}\}_{j=1}^{n_{2}}$. $\pi_{j}^{(2)}$ is considered a neighbour of itself.
Intuitively, if $\pi_{i}^{(1)}$ and $\pi_{j}^{(2)}$ correspond to the same underlying propensity score, their nearest neighbours amongst observations from each modality should be the same. This is measured as a set difference between $\texttt{11}_{ij}$ and $\texttt{12}_{ij}$, and likewise for $\texttt{21}_{ij}$ and $\texttt{22}_{ij}$. Then, a modified Jaccard index is computed as follows. Define
$J_{ij}=|\texttt{11}_{ij}\cap\texttt{12}_{ij}|+|\texttt{21}_{ij}\cap\texttt{22}_{ij}|,$ (14)
the sum of the number of shared neighbours measured in each modality. Then, we compute the following Jaccard distance to populate the unnormalized matching matrix:
$\tilde{M}_{ij}=\frac{J_{ij}}{4k-J_{ij}},$ (15)
where notice that $4k=|\texttt{11}_{ij}|+|\texttt{12}_{ij}|+|\texttt{21}_{ij}|+|\texttt{22}_{ij}|$, since each set contains $k$ distinct neighbours, and thus $0\leq\tilde{M}_{ij}\leq 1$, as with the standard Jaccard index. Then, we normalize each row to produce the final matching matrix:
$M_{ij}=\frac{\tilde{M}_{ij}}{\sum_{i=1}^{n_{1}}\tilde{M}_{ij}}.$ (16)
Note $M_{ij}$ is always well defined because $\pi_{i}^{(1)}$ and $\pi_{j}^{(2)}$ are always considered neighbours of themselves.
Lemma B.1.
$\tilde{M}_{ij}$ has at least one non-zero entry in each of its rows and columns for any number of neighbours $k\geq 1$.
Proof.
We prove that $J_{ij}>0$ for at least one $j$ in each $i$, which is equivalent to $\tilde{M}_{ij}>0$. Fix an arbitrary $i$. $\texttt{21}_{ij}$ by definition is the same set for every $j$. By the assumption of $k\geq 1$ it is non-empty, so there exists $\pi_{j^{*}}^{(2)}\in\texttt{21}_{ij}$. Since $\pi_{j^{*}}^{(2)}$ is a neighbour of itself, we have $\pi_{j^{*}}^{(2)}\in\texttt{22}_{ij^{*}}$, showing that $J_{ij^{*}}>0$. The same reasoning applied to 11 and 12 also shows that $J_{ij}$ for at least one $i$ in each $j$. ∎
Appendix C Omitted Proofs
C.1 Proof of Proposition 3.1
Proof.
Let $x^{(e,t)}$ denote the observed modality and $z^{(t)},u^{(e)}$ be the unique corresponding latent values. By injectivity,
$\pi(x^{(e,t)})$ $=P(t|X^{(e,t)}=x^{(e,t)})$
$=P(t|Z^{(t)}=z^{(t)},U^{(e)}=u^{(e)})$
$=P(t|Z^{(t)}=z^{(t)})=\pi(z^{(t)}),$ (17)
for $e=1,2$, since we assumed $U^{(e)}{\perp\!\!\!\perp}t\mid Z^{(t)}$ in Section 3. Since this holds pointwise, it shows that $\pi(X^{(1,t})=\pi(X^{(2,t)})=\pi(X^{(t)})=\pi(Z^{(t)})$ as random variables. Now, a classical result of Rubin (1974) gives that $Z^{(t)}{\perp\!\!\!\perp}t\mid\pi(Z^{(t)})$, and that for any other function $b$ (a balancing score) such that $Z^{(t)}{\perp\!\!\!\perp}t\mid b(Z^{(t)})$, we have $\pi(Z^{(t)})=g(b(Z^{(t)}))$. The first property written in information theoretic terms yields,
$I(t,Z^{(t)}\mid\pi(Z^{(t)}))=I(t,Z^{(t)}\mid\pi(X^{(t)}))=0,$ (18)
since $\pi(X^{(t)})=\pi(Z^{(t)})$ as random variables, as required. ∎
C.2 Proof of Proposition 3.2
Proof.
In what follows, we write $\pi$ to be the restriction to its domain where it is strictly positive. The $i$-th dimension of the propensity score can be written as
$(\pi(z))_{i}=p(t=i|z)=\frac{p(z|t=i)p(t=i)}{\sum_{i=0}^{T}p(z|t=i)p(t=i)},$ (19)
which, when restricted to be strictly positive, maps to the relative interior of the $T$-dimensional probability simplex. Consider the following transformation:
$h(\pi(z))_{i}=\log\left(\frac{(\pi(z))_{i}}{(\pi(z))_{0}}\right)$ (20)
$=\log(p(z|t=i))-\log(p(z|t=0))+C,$ (21)
where $C=\log(p(t=i))-\log(p(t=0))$ is constant in $z$, and that $h(\pi(z))_{0}\equiv 0$. Ignoring the constant first dimension, we can view $h$ as an invertible map to $\mathbb{R}^{T}$. Under this convention, the map $h\circ\pi:\mathbb{R}^{d}\to\mathbb{R}^{T}$ is smooth ($\log$ is smooth, and the densities are smooth by assumption). Since it is smooth, it cannot be injective if $T<d$ (Spivak, 2018). Finally, since $h$ is bijective, this implies that $\pi$ cannot be injective. ∎
Appendix D Experimental Details
D.1 Evaluation Metrics
We can evaluate how well samples are matched using the ground truth provided by our datasets. In these cases, the dataset sizes are necessarily balanced, so that $n=n_{1}=n_{2}$. In each case, the metric is a function of a $n\times n$ matching matrix $M$, which recall we compute within samples with the same $t$. Our reported results are then averaged over each cluster. We randomize the order of the datasets before performing the matching to avoid pathological cases.
Trace Metric   Assuming the given indices correspond to the true matching, we can compute the average weight on correct matches, which is the normalized trace of $M$:
$\frac{1}{n}\text{Tr}(M)=\frac{1}{n}\sum_{i=1}^{n}M_{ii}.$ (22)
As a baseline, notice that a uniformly random matching that assigns $M_{ij}=1/n$ for each cell yields $\text{Tr}(M)=1$ and hence will obtain a metric of $1/n$. This metric however does not capture potential failure modes of matching. For example, exactly matching one sample, while adversarially matching dissimiliar samples for the remainder also yields a trace of $1/n$, which is equal to that of a random matching.
Latent MSE   Because the image dataset is synthetic, we have access to the ground truth latent values that generated the images, $\mathbf{z}=\{z_{i}\}_{i=1}^{n}$. We compute matched latents as $M\mathbf{z}$, the mean projection according to the matching matrix. Then, to evaluate the quality of the matching, we compute the MSE:
$\text{MSE}(M)=\frac{1}{n}\|\mathbf{z}-M\mathbf{z}\|_{2}^{2}.$ (23)
FOSCTTM   For the CITE-seq dataset, we can use the Fraction Of Samples Closer Than the True Match (FOSCTTM) (Demetci et al., 2022; Liu et al., 2019) as an alternative matching metric. First, we distribute the mass of $\mathbf{x}^{(2)}=\{x_{j}\}_{j=1}^{n}$ to $\mathbf{x}^{(1)}=\{x_{i}\}_{i=1}^{n}$ as $\hat{\mathbf{x}}^{(1)}=M\mathbf{x}^{(2)}$, resulting in a projection of the first modality $\mathbf{x}^{(1)}$ in the space of the second modality. Then, we can compute a cross-modality distance as follows. For each point in $\hat{\mathbf{x}}^{(1)}$, we compute the Euclidean distance to each point in $\mathbf{x}^{(1)}$, and compute the fraction of samples in $\mathbf{x}^{(1)}$ that are closer than the true match. We also repeat this for each point in $\mathbf{x}^{(1)}$, computing the fraction of samples in $\hat{\mathbf{x}}^{(1)}$ in this case. That is, assuming again that the given indices correspond to the true matching, we compute:
$\text{FOSCTTM}(M)=$
$\frac{1}{2n}\bigg{[}\sum_{i=1}^{n}\bigg{(}\frac{1}{n}\sum_{j\neq i}\mathds{1}\{d(\hat{\mathbf{x}}_{i}^{(1)},\mathbf{x}^{(1)}_{j})<d(\hat{\mathbf{x}}_{i}^{(1)},\mathbf{x}^{(1)}_{i})\}\bigg{)}$ (24)
$+\sum_{j=1}^{n}\bigg{(}\frac{1}{n}\sum_{i\neq j}\mathds{1}\{d(\mathbf{x}_{j}^{(1)},\hat{\mathbf{x}}_{i}^{(1)})<d(\mathbf{x}^{(1)}_{j},\hat{\mathbf{x}}_{j}^{(1)})\}\bigg{)}\bigg{]},$ (25)
where notice that this is a function of $M$ through the computation $\hat{\mathbf{x}}^{(1)}=M\mathbf{x}^{(2)}$. As a baseline, we should expect a random matching, when distances between points are randomly distributed, to have an FOSCTTM of $0.5$.
D.2 Models
In this section we describe experimental details pertaining to the propensity score and VAE (Yang et al., 2021). SCOT (Demetci et al., 2022) and scGLUE (Cao & Gao, 2022) are used according to tutorials and recommended default settings by the authors.
Loss Functions   The propensity score approach minimizes the standard cross-entropy loss for both modalities, as implemented in PyTorch 2.0.1. The VAE includes, in addition to the standard ELBO loss (with parameter $\lambda$ on the KL term), two cross-entropy losses based on classifiers from the latent space: one, weighted by a parameter $\alpha$ to classify $t$ as in the propensity score, and another, weighted by a parameter $\beta$, that classifies which modality the latent point belongs to.
Hyperparameters and Optimization   We use the Adam optimizer with learning rate $0.0001$ and one cycle learning rate scheduler. We follow Yang et al. (2021) and set $\alpha=1$, $\beta=0.1$, but found that $\lambda=10^{-9}$ (compared to $\lambda=10^{-7}$ in Yang et al. (2021)) resulted in better performance. We used batch size 256 in both instances and trained for either 100 epochs (image) or 250 epochs (CITE-seq).
Architecture   For the synthetic image dataset, we use an 5-layer convolutional network (channels $=32,54,128,256,512$) with batch normalization and leaky ReLU activations, with linear heads for classification (propensity score and VAE) and posterior mean and variance estimation (VAE). For the VAE, the decoder consists of convolutional transpose layers that reverse those of the encoder. For the CITE-seq dataset, we use a 5-layer MLP with constant hidden dimension $1024$, with batch normalization and ReLU activations (adapted from the fully connected VAE in Yang et al. (2021)) as both the encoder and VAE decoder. We use the same architecture for both modalities, RNA-seq (as we process the top 200 PCs) and protein.
Optimal Transport   We used POT (Flamary et al., 2021) to solve the entropic OT problem, using the log-sinkhorn solver, with regularization strength $\gamma=0.05$.
D.3 Data
Synthetic Data   We follow the data generating process Section 3 to generate coloured scenes of two simple objects (circles, or squares) in various orientations and with various backgrounds. The position of the objects are encoded in the latent variable $z$, which is perturbed by a do-intervention (setting to a fixed value) randomly sampled for each $t$. Each object has an $x$ and $y$ coordinate, leading to a $4$-dimensional $z$, for which we consider $3$ separate interventions each, leading to $12$ different settings. The modality then corresponds to whether the objects are circular or square, and a fixed transformation of $z$, while the modality-specific noise $U$ controls background distortions. Scenes are generated using a rendering engine from PyGame as $f^{(e)}$. Example images are given in Figure 4.
Figure 4: Example pair of synthetic images with the same underlying $z$.
CITE-seq Data
We also use the CITE-seq dataset from Lance et al. (2022) as a real-world benchmark (obtained from GEO accession GSE194122). These consist of paired RNA-seq and surface level protein measurements, and their cell type annotations over $45$ different cell types. We used scanpy, a standard bioinformatics package, to perform PCA dimension reduction on RNA-seq by taking the first 200 principal components. The protein measurements (134-dimensional) was processed in raw form. For more details, see Lance et al. (2022).

### v2

@@
 Abstract
-Multimodal representation learning techniques typically rely on paired samples to learn common representations, but paired samples are challenging to collect in fields such as biology where measurement devices often destroy the samples.
+Multimodal representation learning techniques typically require paired samples to learn shared representations, but collecting paired samples can be challenging in fields like biology, where measurement devices often destroy the samples.
 This paper presents an approach to address the challenge of aligning unpaired samples across disparate modalities in multimodal representation learning.
-We draw an analogy between potential outcomes in causal inference and potential views in multimodal observations, which allows us to use Rubin’s framework to estimate a common space in which to match samples.
-Our approach assumes we collect samples that are experimentally perturbed by treatments, and uses this to estimate a propensity score from each modality, which encapsulates all shared information between a latent state and treatment and can be used to define a distance between samples.
-We experiment with two alignment techniques that leverage this distance—shared nearest neighbours (SNN) and optimal transport (OT) matching—and find that OT matching results in significant improvements over state-of-the-art alignment approaches in both a synthetic multi-modal setting and in real-world data from NeurIPS Multimodal Single-Cell Integration Challenge.
+We draw an analogy between potential outcomes in causal inference and potential views in multimodal observations, allowing us to leverage Rubin’s framework to estimate a common space for matching samples.
+Our approach assumes experimentally perturbed samples by treatments, and uses this to estimate a propensity score from each modality.
+We show that the propensity score encapsulates all shared information between a latent state and treatment, and can be used to define a distance between samples.
+We experiment with two alignment techniques that leverage this distance—shared nearest neighbours (SNN) and optimal transport (OT) matching—and find that OT matching results in significant improvements over state-of-the-art alignment approaches in on synthetic multi-modal tasks, in real-world data from NeurIPS Multimodal Single-Cell Integration Challenge, and on a single cell microscopy to expression prediction task.
 1 Introduction
@@
 However, the effectiveness of multimodal methods hinges on the availability of paired samples—such as images and their associated captions—across data modalities.
-This reliance on paired samples is most obvious in the InfoNCE loss (Gutmann & Hyvärinen, 2010; van den Oord et al., 2018) used in CLIP (Radford et al., 2021) which explicitly learns representations to maximize the true matching between images and their captions.
+This reliance on paired samples is most obvious in the InfoNCE loss (Gutmann and Hyvärinen, 2010; van den Oord et al., 2018) used in CLIP (Radford et al., 2021) which explicitly learns representations to maximize the true matching between images and their captions.
+Figure 1: Visualization of propensity score matching for two modalities (e.g., Microscopy images and RNA expression data).
+We first train classifiers to estimate the propensity score for samples from each modalities; the propensity score reveals the shared information $p(t|z_{i})$, which allows us to re-pair the observed disconnected modalities.
+The matching procedure is then performed within each perturbation class based on the similarity bewteen the propensity scores.
 While paired image captioning data is abundant on the internet, paired multimodal data is often challenging to collect in scientific experiments.
-Take, for instance, experiments in biology, where unpaired data are the norm for technical reasons: RNA sequencing, protein expression assays, and the collection of microscopy images for cell painting experimental assays are all destructive processes.
-Because of this, we cannot collect multiple different measurements from the same cell, and can only explicitly group cells by their experimental condition.
-If we could accurately match unpaired samples across modalities, we could use the aligned samples as proxies for paired samples and apply existing multimodal representation learning techniques.
-However, matching across modalities is challenging for two reasons.
-Firstly, observations from various modalities often exist in entirely different spaces.
-For instance, microscopy images of cells are in pixel space, while gene expression assays provide data in the form of mRNA abundance counts.
-Secondly, even if measurements were taken in the same space, determining an appropriate distance metric for matching remains a non-trivial task.
-For example, when matching images of distinct cells in different microscopy images, we would ideally want a metric that is more sensitive to features describing the underlying state of the cell than features related to biologically-irrelevant cell appearance, such as the orientation of the cell.
-To understand the matching problem more abstractly, we can think of the multiple modalities as potential “views”, $x^{(1)}(z)\in\mathcal{X}^{(1)},x^{(2)}(z)\in\mathcal{X}^{(2)}$ of the same underlying latent state, $z\in\mathcal{Z}$, where we only get to observe one view for any individual unit $i$ (e.g.
-an individual cell in our motivating example).
-Observing, or accurately inferring $z$ would address the first issue by providing a common space $\mathcal{Z}$ in we could match samples.
-However, inferring the latent variable is hopelessly underspecified without making strong assumptions on the data generating processes of both modalities (Xi & Bloem-Reddy, 2023).
-Furthermore, for complex systems such as biological samples, $z$ may still be extremely high-dimensional, and as a result, the second challenge remains: what is the “right” metric is for measuring similarity between samples in this space?
-We address both challenges in this paper by appealing to classical ideas from Rubin (1974), in the case where we additionally observe a label for each unit $t\in\mathcal{T}$, e.g., indexing an experiment that it belongs to.
+For instance, unpaired data are the norm in biology for technical reasons: RNA sequencing, protein expression assays, and the collection of microscopy images for cell painting assays are all destructive processes.
+As such, we cannot collect multiple different measurements from the same cell, and can only explicitly group cells by their experimental condition.
+If we could accurately match unpaired samples across modalities, we could use the aligned samples as proxies for paired samples and apply existing multimodal learning techniques.
+In this paper, we formalize this setting by viewing each modality as a potential measurement, $X^{(1)}(Z)\in\mathcal{X}^{(1)},X^{(2)}(Z)\in\mathcal{X}^{(2)}$, of the same underlying latent state $Z\in\mathcal{Z}$, where we are only able to make a single measurement for each sample unit (e.g.
+an individual cell).
+The task is to reconcile (match) unpaired observations $x^{(1)}$ and $x^{(2)}$ with the same (or maximally similar) $z$.
+Estimating the latent, $Z$, is hopelessly underspecified without making unverifiable assumptions on the system, and furthermore, $Z$ may still be sparse and high-dimensional, leading to inefficient matching.
+This motivates the need for approaches that use only the observable data.
+We identify two major challenges for this problem.
+First, measurements are often made in very different spaces $\mathcal{X}^{(1)}$ and $\mathcal{X}^{(2)}$ (e.g., pixel space and gene expression counts), which make defining a notion of similarity across modalities challenging.
+Second, the measurement process inevitably introduces modality-specific variation that can be impossible to disentangle from the relevant information ($Z$).
+For example in cell imaging, we would not want the matching to depend on irrelevant appearance features such as the orientation of the cell or the lighting of the plate.
+In this paper, we address these challenges by appealing to classical ideas from causal inference (Rubin, 1974), in the case where we additionally observe some label $t$ for each unit, e.g., indexing an experiment.
 By making the assumption that $t$ perturbs the observations via their shared latent state, we identify an observable link between modalities with the same underlying $z$.
-Our key observation is that the propensity score, defined as $p(t|z)$, satisfies three properties: 1) It provides a common space for matching, 2) it maximally coarsens $z$, retaining only relevant information and 3) under certain assumptions, it is estimable from observations of individual modalities alone (Proposition 3.1).
-Figure 1: Our problem setting and proposed solution.
-We assume that two observed modalities (e.g., Microscopy images and RNA expression data) are generated in pairs according to some common latent variable, $z_{i}$, but we only observe the respective modalities separately.
-We further assume the latents are experimentally perturbed by a treatment, $t$, and we get observations from both modalities from a control state and at least one interventional state.
-Under our assumptions, training separate classifiers that predict the intervention label reveals the shared information $p(t|z_{i})$, which allows us to re-pair the observed disconnected modalities.
-In practice, the propensity score is straightforward to estimate: one simply trains two classifiers—one for each modality—to predict which treatment was applied to the respective modalities, and then match across modalities based on the similarity between predicted treatments within each treatment group.
-The proposed methodology is generic across modalities—it can be applied to match observations between any modality for which a classifier can be efficiently trained.
-However, we cannot naively use bipartite matching, as the same sample unit does not appear in both modalities.
-To address this, we use soft matching techniques to estimate the missing modality for each sample unit by combining multiple observations.
-We experiment with two different matching approaches from the recent literature: shared nearest neighbours (SNN) matching (Lance et al., 2022; Cao & Gao, 2022) and optimal transport (OT) matching Villani (2009); Schiebinger et al.
-(2019), and find that OT matching with distances defined on the propensity score lead to significant improvement on real world multi-modal matching tasks from the NeurIPS Multimodal Single-Cell Integration Challenge Lance et al.
-(2022).
-We found similarly large improvements on synthetic image-based benchmarks.
-Finally, we show how our matching procedure can be used for cross-modality prediction tasks.
-We estimate a function that translates between modalities by integrating over the coupling matrix output by our alignment procedure.
-Our approach uses a two-sample stochastic gradient estimator to get an unbiased estimate of the gradient of the resulting loss.
-In our experiments, this approach gave very strong performance, even outperforming methods that had access to the ground truth matching.
-2 Related work
-Unpaired Data   Dealing with unpaired data is important in translation tasks, either for images (Liu et al., 2017; Zhu et al., 2017; Almahairi et al., 2018), or more recently for biological modalities (Amodio & Krishnaswamy, 2018; Yang et al., 2021).
+Under conditions which we discuss in Section 2, the propensity score, defined as $p(t|Z)$, is a transformation of the latent $Z$ that satisfies three remarkable properties (Proposition 3.1): (1) it provides a common space for matching, (2) it is fully identifiable via classification on individual modalities, and (3) it maximally reduces the dimension of $Z$, retaining only the information revealed by the perturbations.
+The practical implementation of the methodology (as illustrated in Fig.
+1) is then straightforward: we train two separate classifiers, one for each modality, to predict the treatment $t$ applied to $X^{(i)}$.
+We then match across modalities based on the similarity between the predicted probabilities (the propensity score) within each treatment group.
+This matching procedure is highly versatile and can be applied to match labeled observations between any modalities for which a classifier can be efficiently trained.
+However, since the same sample unit does not appear in both modalities, we cannot use naive bipartite matching.
+To address this, we use soft matching techniques to estimate the missing modality for each sample unit by allowing matching to multiple observations.
+We experiment with two recent matching approaches: shared nearest neighbours (SNN) matching (Lance et al., 2022; Cao and Gao, 2022) and optimal transport (OT) matching Villani (2009).
+In our experiments, we find that OT matching with distances defined on the proposenity score leads to significant improvement on matching and a downstream cross-modality prediction task on both synthetic and real-world biological data.
+Notably, our prediction method, which leverages the soft matching to optimize an OT projected loss, outperforms supervised learning on the true pairs on CITE-seq data from the NeurIPS Multimodal Single-Cell Integration Challenge (Lance et al., 2022).
+Finally, we applied our method to match single-cell expression data (from a PeturbSeq assay (Dixit et al., 2016)) with single cell crops of image data (Fay et al., 2023).
+We find improved generalization in predicting the distribution of gene expression from the cell imaging data in with unseen perturbations.
+1.1 Related Work
+Unpaired and Multimodal Data
+Learning from unpaired data has long been considered for image translation (Liu et al., 2017; Zhu et al., 2017; Almahairi et al., 2018), and more recently for biological modality translation (Amodio and Krishnaswamy, 2018; Yang et al., 2021).
 In particular, Yang et al.
-(2021) formalize this setting as being generated by a shared underlying latent variable, which has also been identified as a useful modelling assumption for identifiability analyses in multi-view or multi-modal representation learning (Gresele et al., 2020; Sturma et al., 2023).
-Our work also exploits this modelling assumption for theoretical justification, but, unlike previous works, our method does not depend on performing inference over the latent space, thus avoiding the identifiability issue.
-Specifically, since we do not require modality translation, we do not require an encoder-decoder structure (Yang et al., 2021), which embeds modality-specific reconstruction information in the latent space that can be detrimental for the simpler matching task.
-Optimal Transport Matching   Optimal transport is commonly used to solve a similar problem in single-cell biology known as cell trajectory inference.
-In this case, $X^{(1)}$ to $X^{(2)}$ are random variables measured at different time points in a shared (metric) space, and OT matching can be achieved by minimizing the metric (Schiebinger et al., 2019; Tong et al., 2020).
-Recent work (Demetci et al., 2022) tackles our multi-modal alignment problem by using the Gromov-Wasserstein distance, which leverages local metric structure within pairs of points from each modality (Demetci et al., 2022).
-In addition to this “pure” OT approach, Gossi et al.
+(2021) also takes the perspective of a shared latent variable for biological modalities.
+This setting has been studied more generally for multi-view representation learning (Gresele et al., 2020; Sturma et al., 2023) for its identifiability benefits.
+Perturbations and Heterogeneity   Many methods in biology treat observation-level heterogeneity as a nuisance dimension to globally integrate, even when cluster labels are observed (Butler et al., 2018; Korsunsky et al., 2019; Foster et al., 2022).
+This is sensible when clusters correspond to noise rather than the signal of interest.
+However, it is well known in causal representation learning that heterogeneity—particularly heterogeneity arising from perturbations—has theoretical benefits in constraining the solution set (Khemakhem et al., 2020; Squires et al., 2023; Ahuja et al., 2023; Buchholz et al., 2023; von Kügelgen et al., 2023).
+There, the benefits (weakly) increase with the number of perturbations, which is also true of our setting (Proposition 3.2).
+In the context of unpaired data, only Yang et al.
+(2021) explicitly leverage this heterogeneity in their method, while Ryu et al.
+(2024) treat it as a constraint in solving OT.
+Specifically, Yang et al.
+(2021) require their VAE representations to classify experimental labels in addition to reconstructing modalities, while our method is simpler, only requiring the classification objective.
+Notably, Yang et al.
+(2021) treat our objective as a regularizer, but our theory suggests that it is actually primarily responsible for the matching performance.
+Our experiment results coincide with the theoretical insights; requiring reconstruction, as in a VAE, led to worse matching performance with identical model architectures.
+Optimal Transport Matching   OT is a common tool in single-cell biology.
+In cell trajectory inference, the unpaired samples are gene expression values measured at different time points in a shared (metric) space.
+OT matching minimizes this shared metric between time points (Schiebinger et al., 2019; Tong et al., 2020).
+Recent work (Demetci et al., 2022) extends this to our setting where each modality is observed in separate metric spaces by using the Gromov-Wasserstein distance, which computes the difference between the metric evaluated within pairs of points from each modality (Demetci et al., 2022).
+In concurrent work, this approach was recently extended to ensure matching within experimental labels (Ryu et al., 2024).
+In addition to these “pure” OT approaches, Gossi et al.
 (2023) use OT on contrastive learning representations, though this approach requires matched pairs for training, while Cao et al.
 (2022) use OT in the latent space of a multi-modal VAE.
-Perturbations and Heterogeneity   Many methods treat observation-level heterogeneity as another dimension to globally integrate, even when cluster labels are observed (Butler et al., 2018; Korsunsky et al., 2019; Foster et al., 2022).
-This is sensible when clusters correspond to noise rather than the signal of interest.
-However, it is well known in causal representation learning that heterogeneity—particularly heterogeneity arising from different perturbations—has theoretical benefits (Khemakhem et al., 2020; Squires et al., 2023; Ahuja et al., 2023; Buchholz et al., 2023; von Kügelgen et al., 2023).
-There, the benefits (weakly) increase with the number of perturbations, which is also true of our setting (Proposition 3.2).
-In the context of aligning unpaired data, only Yang et al.
-(2021) leverage this heterogeneity in their method.
-Specifically, they require VAE representations to classify experimental labels in addition to reconstructing modalities, while our method only requires a representation to classify labels.
-Yang et al.
-(2021) treat the classifier objective as a regularizer, but our theory suggests that it is responsible for their matching performance.
-In our experiments, we found that requiring reconstruction led to worse matching performance with equivalent architectures.
-3 Propensity Score Multi-modal Matching
-We consider multimodal settings where there exist two potential views, $X^{(e,t)}\in\mathcal{X}^{(e)}$ from two different modalities indexed by $e\in\{1,2\}$ and experiment $t$ that perturbs a shared latent state of these observations.
-The observations will typically be in very different spaces: for example, $\mathcal{X}^{(1,t)}$ may be the space of images of cells under a microscope, and $\mathcal{X}^{(2,t)}$ may be the space of gene expression data.
-This process defines a jointly distributed random variable $(X^{(1,t)},X^{(2,t)},e,t)$, from which we observe only a single modality, its index, and treatment assignment,$\{x_{i}^{(e_{i},t_{i})},e_{i},t_{i}\}_{i=1}^{n}$ (we will always denote random variables by upper-case letters, and samples by their corresponding lower-case letter).
-Our aim is to match or estimate the samples from the missing modality that would correspond to the realization of the missing random variable.
-Importantly, we match within treatment groups, and the treatment variable is only used to learn a common space in which to match.
-We assume each modality arises from a common latent random variable $Z$ as follows,
-$t\sim P_{T},\quad Z^{(t)}\mid t\sim P_{Z}^{(t)},\quad U^{(e)}\sim P_{U}^{(e)},\quad U^{(e)}{\perp\!\!\!\perp}Z,$
+2 Setting
+We consider the setting where there exist two potential views, $X^{(e)}\in\mathcal{X}^{(e)}$ from two different modalities indexed by $e\in\{1,2\}$, and experiment $t$ that perturbs a shared latent state of these observations.
+This defines a jointly distributed random variable $(X^{(1)},X^{(2)},e,t)$, from which we observe only a single modality, its index, and label, $\{x_{i}^{(e_{i})},e_{i},t_{i}\}_{i=1}^{n}$.
+[^1] We will denote random variables by upper-case letters, and samples by their corresponding lower-case letter.
+We aim to match or estimate the samples from the missing modality, which corresponds to the realization of the missing random variable.
+Since $t$ is observed, in practice we match observations within the same label class $t$.
+Formally, we assume each modality is generated by a common latent random variable $Z$ as follows:
+$t\sim P_{T},\ Z^{(t)}\mid t\sim P_{Z}^{(t)},\ U^{(e)}\sim P_{U}^{(e)},\ U^{(e)}{\perp\!\!\!\perp}Z,\ U^{(e)}{\perp\!\!\!\perp}U^{(e^{\prime})},\ X^{(e)}\mid t=f^{(e)}(Z^{(t)},U^{(e)}),$ (1)
+where $t$ indexes the experimental perturbations, and we take $t=0$ to represent a base environment.
+$U^{(e)}$ represents the modality-specific measurement noise that is unperturbed by $t$, and also independent across samples.
+The structural equations $f^{e}$ are deterministic after accounting for the randomness in $Z$ and $U$: it represents the measurement process that captures the latent state.
+For example, in a microscopy image, this would be the microscope and camera that maps a cell to pixels.
+Comparison to Multimodal Generative Models
+Our setting is technically that of a multimodal generative model with latent perturbations.
+However, by focusing on matching rather than generation, we are able to make significantly weaker and more meaningful assumptions while still ensuring the theoretical validity of our method.
+Without the effects of the perturbation, our Eq.
+1 is essentially the same as (Yang et al., 2021, Equation 1) in an abstract sense.
+However, in order to fit the generative model, it is required to formulate explicit models over $f^{(e)}$ and $P_{Z}^{(t)}$, which requires specifying the function class (e.g., continuous) and the space of $Z$ (e.g., $\mathbb{R}^{d}$) as assumptions, even in universal approximation settings.
+In contrast, since we will not directly fit the model Eq.
+1, we do not make any technical assumptions about the generative model.
+Instead, we will make the following assumptions on the underlying data generating process itself.
+Key Assumptions   Our theory makes the following assumptions about the data generating process.
+  • (A1)
 […]
+  • (A2)
+    Injectivity of $f^{(e)}$: $f^{(e)}(z,u)=f^{(e)}(z^{\prime},u^{\prime})\implies(z,u)=(z^{\prime},u^{\prime})$.
+In words, each modality captures enough information to distinguish changes in the underlying state.
+[^2] Note that the injectivity is in the sense of $f$ as a function of both $u$ and $z$, which allows observations that have a shared $z$ but differ by their value in $u$, and the function remains injective.
+For example, rotated images with the exact same content can have a shared $z$, but remain injective due to the rotation being captured in $u$.
 […]
-Under this model, if $Z$ were observable, an optimal matching can be constructed by simply matching the modalities with the most similar $Z$.
-However, $Z$ is latent, and inference on the model described by Section 3 is arguably more difficult than the matching problem itself due to theoretical difficulties such as identifiability and disentangling $Z$ from the modality-specific noise terms $U^{(e)}$.
-Instead, we see that the interventions $t$ provide an observable link between the modalities, thereby revealing information about $Z^{(t)}$.
-Specifically, we use the propensity score with respect to $t$, which we define as
+3 Multimodal Propensity Scores
+Under (1), if $Z$ were observable, an optimal matching can be constructed by simply matching the samples with the most similar $z_{i}$.
+However, the prerequisite of inverting the model and disentangling $Z$ is arguably more difficult than the matching problem itself.
+In particular, $Z$ is unidentifiable without strong assumptions on Eq.
+1 (Xi and Bloem-Reddy, 2023), and even formulating the identifiability problem requires well-specification of the model as a prerequisite.
+We take an alternative approach that is robust to these problems, by using the perturbations $t$ as an observable link to reveal information about $Z$.
+Specifically, we show that the propensity score
 $\pi(z):=P(t|Z=z)\in[0,1]^{T+1},$ (2)
-as a proxy for the latent $Z$.
-Now, although we cannot compute this directly as $Z$ is latent, we make the observation that if $f^{(e)}$ is injective for $e=1,2$, then we can compute the compute the propensity score from each of the observed modalities, since it will be that $\pi(Z^{(t)})=\pi(X^{(e,t)})$, for both $e=1$ and $e=2$.
-Not only does the propensity score reveal shared information, classical causal inference (Rubin, 1974) states that it captures all information shared between the latent and treatment, and does so minimally, in terms of having minimum dimension and entropy.
+is identifiable as a proxy for the latent $Z$ under our assumptions of the data generating process.
+This is a consequence of the injectivity of $f^{(e)}$, since it will be that $\pi(Z)=\pi(X^{(e)})$, $e=1,2$, indicating that we can compute it from either modality.
+Not only does the propensity score reveal shared information, classical causal inference theory (Rubin, 1974) states that it captures all information about $Z$ that is contained in $t$, and does so minimally, in terms of having minimum dimension and entropy.
+Since $t$ contains the only observable information that is useful for matching, the propensity score is hence an optimal compression of the observed information.
 We collect these observations into the following proposition.
 Proposition 3.1.
-In the model described by Section 3, further assume that $f^{(e)}$ are injective for $e=1,2$.
-Then, the propensity score in either modality is equal to the propensity score given by $Z^{(t)}$, i.e., $\pi(X^{(1,t)})=\pi(X^{(2,t)})=\pi(Z^{(t)})$ as random variables.
+In the model described by Eq.
+1, further assume that $f^{(e)}$ are injective for $e=1,2$.
+Then, the propensity scores in either modality is equal to the propensity score given by $Z$, i.e., $\pi(X^{(1)})=\pi(X^{(2)})=\pi(Z)$ as random variables.
 This implies
-$I(t,Z^{(t)}\mid\pi(Z^{(t)}))=I(t,Z^{(t)}\mid\pi(X^{(t)}))=0,$ (3)
-for each $t$, where $I$ is the mutual information. Furthermore, any other function $b(Z^{(t)})$ satisfying $I(t,Z^{(t)}\mid b(Z^{(t)}))=0$ is such that $\pi(Z^{(t)})=f(b(Z^{(t)}))$.
-The proof can be found in the Appendix.
-The above shows that computing the propensity score on either modality is equivalent to computing it on the unobserved shared latent, which captures all the shared information observable in $t$.
-The final statement implies that it is of minimal dimension and entropy, and thus it discards the modality-specific information that may be counterproductive to matching.
-Number of Perturbations   Note that point-wise equality of the propensity score $\pi(z_{1})=\pi(z_{2})$ does not necessarily imply equality of the latents $z_{1}=z_{2}$, due to potential non-injectivity.
-In fact, suppose that $Z^{(t)}$ is supported in $\mathbb{R}^{d}$.
-The propensity score $\pi(z)$ maps into the $T$-dimensional simplex, where $T$ is the number of perturbations, which can be parameterized by coordinates in $\mathbb{R}^{T}$ (essentially, the logits of the classifier).
-Under mild conditions, the propensity score necessarily compresses information about $Z^{(t)}$ if the latent dimension exceeds the number of perturbations, echoing impossibility results from the causal representation learning literature (Squires et al., 2023).
+$I(t,Z\mid\pi(Z))=I(t,Z\mid\pi(X^{(e)}))=0,$ (3)
+for each $e=1,2$, where $I$ is the mutual information. Furthermore, any other function $b(Z)$ satisfying $I(t,Z\mid b(Z))=0$ is such that $\pi(Z)=f(b(Z))$.
+The proof can be found in Appendix C.
+Practically, Proposition 3.1 shows that computing the propensity score on either modality is equivalent to computing it on the unobserved shared latent, which means that it is identifiable, and thus estimable, from the observations alone.
+Furthermore, the estimation does not require modified objectives or architectures for joint multimodal processing, instead they are simple and separate classification problems for each modality.
+Finally, $t$ does not affect $U^{(e)}$ by assumption, and thus the propensity score, being a representation of the information in $t$, discards the modality-specific information that may be counterproductive to matching.
+Therefore, even if $Z$ were observed, it may be sensible to match on its propensity score instead.
+Number of Perturbations   Note that point-wise equality of the propensity score $\pi(z_{1})=\pi(z_{2})$ does not necessarily imply equality of the latents $z_{1}=z_{2}$, due to potential non-injectivity of $\pi$.
+For example, consider $t\in\{0,1\}$, then $\pi(z)$ is a compression to a single dimension $z\to p(t=1\mid z)$.
+Intuitively, collecting data from more perturbations improves the amount of information contained in the label $t$.
+If the latent space is $\mathbb{R}^{d}$, the propensity score necessarily compresses information about $Z^{(t)}$ if the latent dimension exceeds the number of perturbations, echoing impossibility results from the causal representation learning literature (Squires et al., 2023).
 Proposition 3.2.
-Suppose that $P_{Z}^{(t)}$ has a smooth density $p(z|t)$ for each $t=0,\dots,T$. Then, if $T<d$, the propensity score $\pi$, restricted to its strictly positive part, is non-injective.
-The proof can be found in the Appendix.
+Let $Z^{(t)}\in\mathbb{R}^{d}$.
+Suppose that $P_{Z}^{(t)}$ has a smooth density $p(z|t)$ for each $t=0,\dots,T$.
+Then, if $T<d$, the propensity score $\pi$, restricted to its strictly positive part, is non-injective.
+The proof can be found in Appendix C.
 Note the above only states an impossibility result when $T<d$.
-When $T\geq d$, it can be seen from the proof of Proposition 3.2 that the injectivity of the propensity score depends on the injectivity of the following expression in $z$:
-$f(z)=\begin{bmatrix}\log(p(z|t=1))-\log(p(z|t=0))\\ \vdots\\ \log(p(z|t=T))-\log(p(z|t=0))\end{bmatrix},$ (4)
+More generally, it can be seen from the proof of Proposition 3.2 that the injectivity of the propensity score depends on the injectivity of the following expression in $z$:
+$g(z)=\begin{bmatrix}\log(p(z|t=1))-\log(p(z|t=0))\\ \vdots\\ \log(p(z|t=T))-\log(p(z|t=0))\end{bmatrix},$ (4)
 which then depends on the latent process itself.
-If the above is non-injective, this is a fundamental indeterminacy that cannot be resolved without making strong assumptions on point-wise latent variable recovery, as we have already seen that the propensity score contains the maximal shared information, as shown in Proposition 3.1.
-Nonetheless, note that $f$ in Equation 4 is injective if any of the subset of its entries are, clearly outlining the benefits of collecting data from a larger number of perturbations for matching.
-Key Assumptions   Our results thus far rely on two key modelling assumptions,
-  • (A1): $t$ has a non-trivial effect on $Z$, but does not affect $u^{(e)}$, implying that interventions are able to target the common underlying process without changing modality-specific properties.
-  • (A2): Injectivity of $f^{(e)}$ in the propensity score derivation, $f^{(e)}(z,u)=f^{(e)}(z^{\prime},u^{\prime})\Rightarrow(z,u)=(z^{\prime},u^{\prime})$, which implies each modality contains complete information about the underlying process.
-(A1) is justifiable insofar as modalities represent different measurements of an isolated system, such as in biological studies where $Z$ might correspond to an underlying cell state and modalities refer to different single-cell measurements.
-However, in practice, even different measurement devices may be more or less sensitive to the biological variation implied by $t$.
-We weaken this assumption in a simplified case in Section 3.
-(A2) requires that no information about the latent state is lost in the observation process.
-Note that the injectivity is in the sense of $f$ as a function of both $u$ and $z$, which allows observations that have a shared $z$ but differ by their value in $u$, and the function remains injective.
-For example, rotated images with the exact same content can have a shared $z$, but remain injective due to the rotation being captured in $u$.
+If the above mapping is non-injective, this represents a fundamental indeterminacy that cannot be resolved without making strong assumptions on point-wise latent variable recovery.
+As we have already established in Proposition 3.1, the propensity score contains the maximal shared information across modalities.
+Nonetheless, collecting data form a larger number of perturbations is clearly beneficial for matching, since $g$ in Eq.
+4 is injective if any of the subset of its entries are.
+4 Estimation and Matching
+For the remainder of the paper, we drop the notation $e$ and use $(x_{i},t_{i})$ to denote observations from modality 1, and $(x_{j},t_{j})$ to denote observations from modality 2.
+Given a multimodal dataset with observations $\{(x_{i},t_{i})\}_{i=1}^{n_{1}}$ and $\{(x_{j},t_{j})\}_{j=1}^{n_{2}}$, we wish to compute a matching matrix (or coupling) between the two modalities.
+We define a $n_{1}\times n_{2}$ matching matrix $M$ where $M_{ij}$ represents the likelihood of $x_{i}$ being matched to $x_{j}$.
+Since $t$ is observed, we always perform matching only within observations with the same value of $t$, so that in practice we obtain a matrix $M_{t}$ for each $t$.
+Our method approximates the propensity scores by training separate classifiers that predicts $t$ given $x$ for each modality.
+We denote the estimated propensity score by $\pi_{i}$ and $\pi_{j}$ respectively, where
+$\pi_{i}\approx\pi(x_{i})=P(T=t\mid X_{i}^{(e)}=x_{i}).$ (5)
+This yields the transformed datasets $\{\pi_{i}\}_{i=1}^{n_{1}}$ and $\{\pi_{j}\}_{j=1}^{n_{2}}$, where $\pi_{i}$, $\pi_{j}$ are in the $T$ dimensional simplex.
+We use this correspondence to compute a cross-modality distance function:
+$d(x_{i},x_{j}):=d^{\prime}(\pi_{i},\pi_{j}).$ (6)
+In practice, we typically compute the Euclidean distance in $\mathbb{R}^{T}$ of the logit-transformed classification scores, but any metric over a bijective transformation of the propensity scores are also theoretically valid.
+Given this distance function, we use existing matching techniques to constructing a matching matrix.
+In our experiments, we found that OT matching gave the best performance, but we also evaluated Shared Nearest Neighbour matching; details of the latter can be found in Appendix B.
+Optimal Transport Matching
+The propensity score distance allows us to easily compute a cost function associated with transporting mass between modalities, $c(x_{i},x_{j})=d^{\prime}(\pi_{i},\pi_{j})$.
+Let $p_{1},p_{2}$ denote the uniform distribution over $\{\pi_{i}\}_{i=1}^{n_{1}}$ and $\{\pi_{j}\}_{j=1}^{n_{2}}$ respectively.
+Discrete OT aims to solve the problem of optimally redistributing mass from $p_{1}$ to $p_{2}$ in terms of incurring the lowest cost.
+Let $C_{ij}=c(x_{i},x_{j})$ denote the $n_{1}\times n_{2}$ cost matrix.
+The Kantorovich formulation of optimal transport aims to solve the following constrained optimization problem:
+$\min_{M}\sum_{i}^{n_{1}}\sum_{j}^{n_{2}}C_{ij}M_{ij},\quad M_{ij}\geq 0,\quad M\mathbf{1}=p_{1},\quad M^{\top}\mathbf{1}=p_{2}.$ (7)
+This is a linear program, and for $n_{1}=n_{2}$, it can be shown that the optimal solution is a bipartite matching between $\{\pi_{i}\}_{i=1}^{n_{1}}$ and $\{\pi_{j}\}_{j=1}^{n_{2}}$.
+We refer to this as exact OT; in practice we add an entropic regularization term, resulting in a soft matching, that ensures smoothness and uniqueness, and can be solved efficiently using Sinkhorn’s algorithm.
+Entropic OT takes the following form:
+$\min_{M}\sum_{i}^{n_{1}}\sum_{j}^{n_{2}}C_{ij}M_{ij}-\lambda H(M),\quad M_{ij}\geq 0,\quad M\mathbf{1}=p_{1},\quad M^{\top}\mathbf{1}=p_{2},$ (8)
+where $H(M)=-\sum_{i,j}M_{ij}\log(M_{ij})$, the entropy of the joint distribution implied by $M$.
+This approach regularizes towards a higher entropy solution, which has been shown to have statistical benefits (Genevay et al., 2018), but nonetheless for small enough $\lambda$ serves as a computationally appealing approximation to exact OT.
+5 Downstream Tasks
+The matching matrix $M$ can be seen as defining an empirical joint distribution over the samples in each modality.
+The OT approach in particular makes this explicit.
+Each row is proportional to the probability that each sample $i$ from modality (1) is matched to sample $j$ in modality (2), i.e., $M_{i,j}=P(x_{j}|x_{i})$.
+We can thus use $M$ to obtain pseudosamples for any learning task that uses paired samples by $(x_{i},\hat{x}_{j})$, where $\hat{x}_{j}$ is obtained by sampling from the conditional distribution defined by $M$, or by a suitable conditional expectation, e.g., the barycentric projection (conditional mean) as $E_{M}\left[X_{j}\mid X_{i}=x_{i}\right]=\sum_{j}M_{i,j}x_{j}$.
+In what follows, we describe a cross-modality prediction method based on both barycentric projection and stochastic gradients according to $M_{i,j}$.
+Cross-modality prediction
+We can use the matching matrix to design a method for cross-modality prediction/translation.
+The following MSE loss corresponds to constructing a prediction function $f_{\theta}$ such that the barycentric projection $E_{M}\left[f_{\theta}(X_{j})\mid X_{i}=x_{i}\right]$, under $M$ minimizes the squared error for predicting $x_{i}$:
+$\mathcal{L}(\theta):=\sum_{i}(x_{i}-\sum_{j}M_{i,j}f_{\theta}(x_{j}))^{2}.$ (9)
+However, this requires evaluating $f_{\theta}$ for all $n_{2}$ examples from modality (2) for each of the $n_{1}$ examples in modality (1).
+In practice, we can avoid this cost with stochastic gradient descent by sampling from modality $(2)$ via $M_{i\cdot}$ for each training example $(1)$.
+To obtain an unbiased estimate of $\nabla_{\theta}\mathcal{L}$, we need two independent samples from modality (2) for each sample from modality (1),
+$\nabla\mathcal{L}(\theta)\approx$ $-2\left(x_{i}-f_{\theta}(\dot{x}_{j})\right)\nabla_{\theta}f_{\theta}(\ddot{x}_{j})\quad\dot{x}_{j},\ddot{x}_{j}\sim P(x_{j}|x_{i}).$ (10)
+By taking two samples as in Eq.
+10, we get an unbiased estimator of $\nabla\mathcal{L}(\theta)$, whereas a single sample would have resulted in optimizing an upper-bound on equation (9); for details, see Hartford et al.
+(2017) where a similar issue arises in the gradient of their causal effect estimator.
+We thus refer to prediction models trained via Eq.
+10 as unbiased.
+6 Experiments
+We present a comprehensive evaluation of our proposed methodology on three distinct datasets: (1) synthetic paired images, (2) single-cell CITE-seq dataset (simultaneous measurement of single-cell RNA-seq and surface protein measurements) (Stoeckius et al., 2017), and (3) Perturb-seq and single-cell image data.
+In the first two cases, there is a ground-truth matching that we use for evaluation, but samples are randomly permuted during training.
+This allows us to exactly compute the quality of the matching in comparison to the ground truth.
+The final dataset is a more realistic setting where ground truth paired samples do not exist, and matching becomes necessary in practice.
+In this case, we compute distributional metrics to compare our proposed methodology against other baselines.
+Experimental Details
+All models for the experiments are implemented using Torch v2.2.2 (Paszke et al., 2017) and Pytorch Lightning v2.2.4 (Falcon and PyTorch Lightning Team, 2023).
+The classifier used to estimate the propensity score is always a linear head on top of an encoder $E_{i}$, which is specific to each modality and dataset.
+All models are saved at the optimal validation loss to perform subsequent matching.
+Shared nearest neighbours (SNN) is implemented using scikit-learn v1.4.0 (Pedregosa et al., 2011) using a single neighbour, and OT is implemented using the Sinkhorn algorithm as implemented in the pot v0.9.3 package (Flamary et al., 2021).
+Both SNN and OT use the Euclidean distance as the metric.
+Whenever random variation can affect the results of the experiments, we report quantiles corresponding to variation from different random seeds.
+Additional experimental details are provided in Appendix D.
+Description of Baselines
+Our main baseline, which we evaluate against on all three datasets, is matching using representations learned by the multimodal VAE of Yang et al.
+(2021), which is the only published method that is able to leverage perturbation labels for unpaired multimodal data (they refer to the labels as “prior information”).
+The standard multimodal VAE loss is a reconstruction loss based on encoder and decoders $E_{i}$, $D_{i}$ for each modality, plus a latent invariance loss that aims to align the modalities in the latent space.
+In our setting, the multimodal VAE loss further includes an additional label classification loss from the latent space of each modality, i.e., encouraging the encoder to simultaneously learn $P(t\mid E_{i}(x_{i}))$.
+This additional objective, which acts as a regularizer for the multimodal VAE, is exactly the loss for our proposed method.
+To ensure a fair comparison, we always use the same architecture in the encoders $E_{i}$ of multimodal VAE and in our propensity score classifier.
+The performance differences between propensity score matching and multimodal VAE then represent the effects of the VAE reconstruction objective and latent invariance objectives.
+For additional baselines, we also compare against a random matching, where the samples are matched with equal weight within each perturbation as a sanity check.
+For datasets (1) and (2), we also compare against Gromov-Wasserstein OT (SCOT) (Demetci et al., 2022) computed separately within each perturbation.
+SCOT uses OT directly by computing a cost function derived based on pairwise distances within each modality, thus learning a local description of the geometry which can be compared between modalities.
+For the CITE-seq dataset, we also compare against matching using a graph-linked VAE, scGLUE (Cao and Gao, 2022), where the graph is constructed from linking genes with the associated proteins.
+Evaluation Metrics
+We use the known ground truth matching to compute performance metrics on datasets (1) and (2).
+The trace and FOSCTTM (Liu et al., 2019) measure how much weight $M$ places on the true pairing.
+However, this is not necessarily indicative of downstream performance as similar, but not exact matches are penalized equally to wildly incorrect matches.
+For this reason, we also measure the latent MSE for dataset (1) and the performance of a CITE-seq gene–to–protein predictive model based on the learned matching for dataset (2).
+For more details, see Section D.1.
+6.1 Experiment 1: Synthetic Interventional Images
+Data   We followed the data generating process Eq.
+1 with a latent variable $Z$ encoding the coordinates of two objects.
+Perturbations represent different do-interventions on the different dimensons of $Z$.
+The difference between modalities corresponds to whether the objects are circular or square, and a fixed transformation of $Z$, while the modality-specific noise $U$ controls background distortions.
+Model and Evaluation   We used a convolutional neural network adapted from Yang et al.
+(2021) as the encoder.
+We report two evaluation metrics: (1) the trace metric, and (2) the MSE between the matched and the true latents.
+The latent MSE metric does not penalize close neighbours of the true match (i.e.
+examples for which $\|z_{i}-z_{i}^{*}\|$ is small) as heavily as the trace metric.
+These “near matches” will typically still be useful on downstream multimodal tasks.
+Results   In Table 1, metrics are computed on a held out test set over 12 groups corresponding to interventions on the latent position, with approximately 1700 observations per group.
+A random matching, with weight $1/n$, will hence have a trace metric of of $1/1700\approx 0.588\times 10^{-3}$.
+This implies, for example, that the median performance of PS+OT is approximately 31 times that of random matching.
+On both metrics, we found that propensity scores matched with OT (PS + OT) consistently outperformed other matching methods on both metrics.
+Synthetic Image Data CITE-seq Data
+Method MSE ($\downarrow$) Trace ($\uparrow$) FOSCTTM ($\downarrow$) Trace ($\uparrow$)
+Med (Q1, Q3) Med (Q1, Q3) ×$\times 10^{-3}$ Med (Q1, Q3) Med (Q1, Q3)
+PS+OT 0.0316 18.329 0.3049 0.1163
+(0.0300, 0.0330) (17.068, 18.987) (0.3008, 0.3078) (0.1093, 0.1250)
+VAE+OT 0.0324 7.733 0.3953 0.0814
+(0.0316, 0.0350) (7.473, 7.794) (0.3912, 0.4045) (0.0777, 0.8895)
+PS+SNN 0.0552 7.924 0.3126 0.0941
+(0.0530, 0.0558) (7.569, 9.504) (0.3121, 0.3160) (0.0880, 0.0989)
+VAE+SNN 0.0622 3.116 0.3816 0.0612
+(0.0571, 0.0676) (2.818, 3.213) (0.3760, 0.3822) (0.0588, 0.0634)
+SCOT 0.0354 0.5964 0.4596 0.0200
+GLUE+SNN - - 0.4412 0.0362
+GLUE+OT - - 0.5309 0.0323
+Random 0.0709 - - -
+(0.0707, 0.0714)
+Table 1: Alignment metrics results using synthetic interventional image dataset and CITE-seq data.
+6.2 Experiment 2: CITE-Seq Data
+Data   We used the CITE-seq dataset from the NeurIPS 2021 Multimodal single-cell data integration competition (Lance et al., 2022), consisting of paired RNA-seq and surface level protein measurements over $45$ cell types.
+In the absence of perturbations, we used the cell type as the observed label to classify and match within.
+Note the cell types are determined by consensus by pooling annotations from marker genes/proteins.
+In most cells, the annotations from each modality agreed, suggesting that the label is independent from the modality-specific noise.
+We used the first 200 principal components as the gene expression modality, and normalized (but otherwise raw) protein measurements as input.
+Model and Evaluation   We used fully-connected MLPs as encoders.
+To assess matching, we report (1) the trace, and (2) the Fraction Of Samples Closer Than the True Match (FOSCTTM) ((Demetci et al., 2022), (Liu et al., 2019)) (lower is better, 0.5 corresponds to random guessing).
+To evaluate against a downstream task, we also compared the performance of random and VAE matching procedures, as well as directly using the ground truth ($M_{ii}=1$), on predicting protein levels from gene expression.
+We trained a 2-layer MLP (the same architecture for all matchings) with both MSE loss and the unbiased procedure as described in Section 5 using pseudosamples sampled according to the matching matrix.
+We evaluated the predictive models against ground truth pairs by computing the prediction $R^{2}$ (higher is better) on a held-out, unpermuted, test set.
+Results   In Table 1, metrics are computed on a held-out test set averaged over 45 cell types with varying observation counts per group.
+While interpreting the average trace can be challenging due to group size variations, OT matching on PS consistently outperformed other methods both within and across groups.
+In these experiments, OT matching on PS was consistently the top performer, often followed by SNN matching on PS or OT matching on VAE embeddings.
+We present downstream task performance in Table 2.
+Note that $R^{2}$ is computed using the sample average across possibly multiple cell types, which explains why random matching within each cell type results in non-zero $R^{2}$ (see Section D.1).
+We found that PS + OT matching outperforms other methods on this task.
+Surprisingly, the PS + OT prediction model performed even better on average than training with the standard MSE loss on ground truth pairings (though confidence intervals overlap).
+This highlights the potential benefit of soft (OT) matching as a regularizer, beyond that of simply reconciling most likely pairs: the soft matching effectively averages over modality specific variation from samples with similar latent states in a manner analogous to data augmentation (with an unknown group action).
+6.3 Experiment 3: PerturbSeq and Single Cell Images
+CITE-seq Data PerturbSeq/Single Cell Image Data
+Method MSE Loss Unbiased Loss In Distribution Out of Distribution
+$R^{2}$ Med (Q1, Q3) ($\uparrow$) $R^{2}$ Med (Q1, Q3) ($\uparrow$) KL Med (Q1, Q3) ($\downarrow$) KL ($\downarrow$)
+Random 0.138 0.173 58.806 51.310
+(0.137, 0.140) (0.170, 0.173) (58.771, 60.531)
+VAE+OT 0.149 0.114 55.483 47.910
+(0.118, 0.172) (0.079, 0.159) (55.410, 56.994)
+PS+OT 0.217 0.233 50.967 43.554
+(0.206, 0.223) (0.207, 0.250) (50.898, 52.457)
+True Pairs 0.224 - - -
+(0.223, 0.226)
+Table 2: Cross-modal prediction results using CITE-seq data and PerturbSeq/single cell image data including an out of distribution distance evaluation for PerturbSeq/single cell images.
+Data   We collected PerturbSeq data (200 genes) and single-cell images of HUVEC cells with 24 gene perturbations and a control perturbation, resulting in 25 total labels across both modalities.
+As preprocessing, we embed the raw PerturbSeq counts into a 128-dimensional space using scVI (Lopez et al., 2018) and the cell images into a 1024-dimensional space using a pre-trained Masked Autoencoder (He et al., 2022; Kraus et al., 2023) to train our gene expression and image classifiers.
+Model and Evaluation   We used a fully connected 2-layer MLP as the encoder for both PerturbSeq and cell image classifiers.
+Similarly to the CITE-seq dataset, we evaluated the matchings based on downstream prediction of gene expression from (embeddings of) images.
+We used the unbiased procedure to minimize the projected loss Eq.
+9 and evaluated on two held-out sets, one consisting of in-distribution samples from the 25 perturbations the classifier was trained on, and an out-of-distribution set consisting of an extra perturbation not seen in training.
+In the absence of ground truth matching, we assessed three distributional metrics between the actual and predicted gene expression values within each perturbation: the L2 norm of the difference in means, the Kullback-Leibler (KL) divergence, and 1-Wasserstein distance (lower indicates better alignment).
+We report inverse cell-count weighted averages over each perturbation group.
+Each metric measures a slightly different aspect of fit—the L2 norm reports a first-order deviation, while the KL divergence is an empirical estimate of the deviation of the underlying predicted distribution, while the 1-Wasserstein distance measures deviations in terms of the empirical samples themselves.
+Note that matching is performed using classifiers trained on scVI embeddings, but the cross-modal predictions are generated in the original log transformed gene expression space (i.e.
+we predicted actual observations, not embeddings).
+We also evaluated distance measures on an out-of-distribution gene perturbation that was not used in either the matching or training of the translation model.
+Results   We present KL divergence values for in-distribution and out-of-distribution in Table 2.333We computation of in-distribution metrics using random subsamples from the test set.
+The out-of-distribution metric was computed on a small dataset with a single perturbation and subsamples were not needed.
+Additional metrics show similar patterns and can be found in Section D.4.
+OT + PS matching consistently outperforms its VAE counterpart both on in-distribution and out-of-distribution metrics, supporting our findings on the CITE-seq data to the case where ground truth pairs are not available.
+6.4 Validation Monitor
+As in our Perturb-seq and cell imaging example, the ground truth matching is typically unknown in real problems.
+It is hence desirable to have an observable proxy of the matching performance as a validation during hyperparameter tuning.
+Figure 2 demonstrates that the propensity score validation loss (cross-entropy) empirically satisfies this role in our CITE-seq experiments, where lower validation loss corresponds to better matching performance, as if it were computed with the ground truth.
+By contrast, we found that the optimal VAE, in terms of matching, had higher validation loss.
+This empirically supports our intuition that the reconstruction loss minimization requires the VAE to capture modality specific information, i.e., the $U^{(e)}$ variables, which hinders its matching performance.
+Figure 2: VAE and classifier validation metrics on the CITE-seq dataset.
+Notice that validation cross-entropy inversely tracks the ground truth matching metrics, and thus can be used as a proxy in practical settings where the ground truth is unknown.
+The same pattern does not hold for the VAE (Yang et al., 2021), which we suspect is because reconstruction is largely irrelevant for matching.
+7 Limitations
+Our methods are limited to settings where we have some signal to play the role of an experiment label, but we believe this is where these methods are most needed.
+Matching is impossible in general—e.g., if you tried to match modalities that have no shared information, it would clearly fail—but our theory formally articulates both where we expect this method to succeed and its limitations.
+Both (A1) and (A2) are strong assumptions, but the empirical results suggest the method is fairly robust to failures.
+8 Conclusion
+This work presents a simple algorithm for aligning unpaired data from different modalities using propensity scores.
+The method is very general, requiring only a classifier to be trained on each modality, and demonstrates excellent matching performance, which we validate both theoretically and empirically.
+We also showcase the effectiveness of the matching algorithm in a downstream cross-modality prediction task, achieving better generalization compared to random matching, VAE-based matching, and even the ground truth matching on the evaluated dataset.
+This improved generalization over the ground truth may be attributed to implicitly enforcing invariance to modality-specific information; a rigorous investigation of this phenomenon would be interesting for further investigation.
+9 Acknowledgements
+We are extremely grateful for the discussions with many external collaborators and colleagues at Recursion that lead to this work.
+The original ideas for this work stemed from conversations with Alex Tong with feedback from Yoshua Bengio at Mila.
+We received a lot of helpful feedback from all of our colleagues at Valence Labs, especially Berton Earnshaw and Ali Denton.
+The single cell image experiments are built on code originally written by Oren Kraus and his team.
+Appendix A Relaxing (A1)
 Relaxing Assumption 1
@@
 where we do not necessarily require $U^{(e)}{\perp\!\!\!\perp}t\mid Z^{(t)}$, and thus we obtain
-$\pi(x^{(1,t)})=P(t|Z^{(t)}=z^{(t)},U^{(1)}=u^{(1)})\neq$ (6)
-$P(t|Z^{(t)}=z^{(t)},U^{(2)}=u^{(2)})=\pi(x^{(2,t)}),$
+$\pi(x^{(1,t)})=P(t|Z^{(t)}=z^{(t)},U^{(1)}=u^{(1)})\neq P(t|Z^{(t)}=z^{(t)},U^{(2)}=u^{(2)})=\pi(x^{(2,t)}),$ (12)
 see the proof of Proposition 3.1 for details.
@@
 Then, exact OT will still recover the ground truth matching.
-See Figure 2 for a visual example of this type of monotonicity.
+See Fig.
+3 for a visual example of this type of monotonicity.
 For example, suppose that $t$ is a chemical perturbation of a cell, and thus $\pi_{i}^{(1)}$, $\pi_{j}^{(2)}$ can be seen as a measure of biological response to the perturbation, e.g., in a treated population, $\pi_{i_{1}}>\pi_{i_{2}}$ indicates samples $i_{1}$ had a stronger response than sample ${i_{2}}$, as perceived by the first modality indexed by $i$.
 Then, this monotonocity states that we should see the same $\pi_{j_{1}}>\pi_{j_{2}}$ in the other modality as well, if the samples $i_{1}$ and $i_{2}$ truly corresponded to $j_{1}$ and $j_{2}$.
-Figure 2: OT matching allows for $t$ to have different effects on the modality specific information, here $u_{i}^{(1)}$ and $u_{i}^{(2)}$, as long as they can be written as transformations that preserve the relative order within modalities.
-Exact OT in 1-d always matches according to the relative ordering, and thus exhibits this type of “no crossing” behaviour shown in the figure on the left.
-The figure on the right shows a case where we would fail to correctly match across modalities because of the crossing shown in orange.
-When $t$ is not a binary treatment and the propensity scores are multidimensional, a condition known as cyclic monotonicity is sufficient for recovering the true matching, see Appendix A for details.
-However, unlike the 1-d case, we are not aware of more intuitive conditions under which we should expect cyclic monotonicity to hold.
-4 Methodology
-For the remainder, we will drop $(e,t)$ from our notation, and use $(x_{i},t_{i})$ to denote observations from modality 1, and $(x_{j},t_{j})$ to denote observations from modality 2 whenever it does not cause confusion to do so.
-Given a multimodal dataset with observations $\{(x_{i},t_{i})\}_{i=1}^{n_{1}}$ and $\{(x_{j},t_{j})\}_{j=1}^{n_{2}}$, we wish to compute a matching matrix (or coupling) between the two modalities.
-We define a $n_{1}\times n_{2}$ matching matrix $M$ where $M_{ij}$ represents the likelihood of $x_{i}$ being matched to $x_{j}$.
-We always perform matching only within observations with the same value of $t$, so that in practice we obtain a matrix $M_{t}$ for each $t$.
-Our method approximates the propensity scores by training separate classifiers that predicts $t$ given $x$ for each modality.
-We denote the estimated propensity score by $\pi_{i}$ and $\pi_{j}$ respectively, where
-$\pi_{i}\approx\pi(x_{i})=P(T=t\mid X_{i}^{(e,t)}=x_{i}).$ (9)
-This results in the transformed datasets $\{\pi_{i}\}_{i=1}^{n_{1}}$ and $\{\pi_{j}\}_{j=1}^{n_{2}}$, where $\pi_{i}$, $\pi_{j}$ are in the $T$ dimensional simplex.
-We use this correspondence to compute a cross-modality distance function:
-$d(x_{i},x_{j}):=d^{\prime}(\pi_{i},\pi_{j}),$ (10)
-where in practice, we typically compute the Euclidean distance in $\mathbb{R}^{T}$ of the logit-transformed classification scores.
-Given this distance function, we rely on existing matching techniques to constructing a matching matrix.
-In our experiments, we found OT matching gave the best performance, but we also evaluated Shared Nearest Neighbour matching; details of the latter can be found in Appendix B.
-4.1 Optimal Transport Matching
-The propensity score distance allows us to easily compute a cost function associated with transporting mass between modalities, $c(x_{i},x_{j})=d^{\prime}(\pi_{i}\pi_{j})$.
-Let $p_{1},p_{2}$ denote the uniform distribution over $\{\pi_{i}\}_{i=1}^{n_{1}}$ and $\{\pi_{j}\}_{j=1}^{n_{2}}$ respectively.
-Discrete OT aims to solve the problem of optimally redistributing mass from $p_{1}$ to $p_{2}$ in terms of incurring the lowest cost.
-Let $C_{ij}=c(x_{i},x_{j})$ denote the $n_{1}\times n_{2}$ cost matrix.
-The Kantorovich formulation of optimal transport aims to solve the following constrained optimization problem:
-$\min_{M}\sum_{i}^{n_{1}}\sum_{j}^{n_{2}}C_{ij}M_{ij},\quad M_{ij}\geq 0,\quad M\mathbf{1}=p_{1},\quad M^{\top}\mathbf{1}=p_{2}.$
-This is a linear program, and for $n_{1}=n_{2}$, it can be shown that the optimal solution is a bipartite matching between $\{\pi_{i}\}_{i=1}^{n_{1}}$ and $\{\pi_{j}\}_{j=1}^{n_{2}}$.
-We refer to this as exact OT; in practice we add an entropic regularization term, resulting in a soft matching, that ensures smoothness and uniqueness, and can be solved efficiently using Sinkhorn’s algorithm.
-Entropic OT takes the following form:
-$\min_{M}\sum_{i}^{n_{1}}\sum_{j}^{n_{2}}C_{ij}M_{ij}-\lambda H(M),\quad M_{ij}\geq 0,\quad M\mathbf{1}=p_{1},\quad M^{\top}\mathbf{1}=p_{2},$
-where $H(M)=-\sum_{i,j}M_{ij}\log(M_{ij})$, the entropy of the joint distribution implied by $M$.
-This approach regularizes towards a higher entropy solution, which has been shown to have statistical benefits (Genevay et al., 2018), but nonetheless for small enough $\lambda$ serves as a computationally appealing approximation to exact OT.
-5 Cross-modality prediction
-Given a matching matrix $M$, we can interpret this as the probability that each sample, $i$, from modality (1) is matched to sample $j$ in modality (2); that is $M_{i,j}=P(x^{(2)}_{j}|x^{(1)}_{i})$.
-As a downstream task, and as an evaluation metric for $M$, we can use this matching to estimate a cross-modal prediction model, $f_{\theta}$, that maps from modality (2) to (1) by minimizing the following loss,
-$\mathcal{L}(\theta):=\sum_{i}(x_{i}^{(1)}-M_{i}f_{\theta}(x_{j}^{(2)}))^{2}.$ (11)
-However, this requires evaluating $f_{\theta}$ for all $n_{2}$ examples from modality (2) for each of the $n_{1}$ examples in modality (1).
-Of course, we can avoid this cost with stochastic gradient descent by sampling from modality $(2)$ via $M_{i\cdot}$ for each training example $(1)$, but to get an unbiased estimate of $\nabla_{\theta}\mathcal{L}$, we need two samples from modality (2) for each sample from modality (1),
-$\nabla\mathcal{L}(\theta)\approx$ $-2\left(x_{i}^{(1)}-f_{\theta}(\dot{x}_{j}^{(2)})\right)\nabla_{\theta}f_{\theta}(\ddot{x}_{j}^{(2)})\quad$ $\dot{x}_{j}^{(2)},\ddot{x}_{j}^{(2)}\sim P(x^{(2)}_{j}|x^{(1)}_{i}).$ (12)
-By taking two samples as in equation (12), we get an unbiased estimator of $\nabla\mathcal{L}(\theta)$, whereas a single sample would have resulted in optimizing an upper-bound on equation (11); for details, see Hartford et al.
-(2017) where a similar issue arises in the gradient of their causal effect estimator.
-6 Experiments
-We evaluate our proposed methodology on two datasets.
-The first is a synthetic interventional image dataset generated satisfying the assumptions of Section 3.
-The second is a real-world single-cell CITE-seq data obtained from the NeurIPS 2021 Multimodal single-cell data integration competition (Lance et al., 2022), which provides a ground truth matching by allowing for a small number of cell surface proteins to be measured simultaneously to RNA sequencing.
-Note the CITE-seq dataset is not interventional—we use the cell type as the classification target $t$ instead.
-All details are made available in the Appendix.
-In both cases the ground truth matching is known to make evaluation possible, but hidden during training.
-Baselines   In terms of the problem setting, the most closely related method is Gromov-Wasserstein optimal transport (SCOT) (Demetci et al., 2022), which explicitly computes a cross-modality cost for optimal transport matching by leveraging the local geometry within each modality.
-For the CITE-seq data, graph-linked VAE (scGLUE), which has access to specific biological metadata that connects the biological modalities, can also be used for matching.
-[^2] A variant of scGLUE was the top entrant in the NeurIPS 2021 Multi-modal single-cell data integration competition, from which we source our data.
-However, ground truth matchings were made available in the training set in that case, which fundamentally changes the nature and difficulty of the problem.
-However, neither of the preceding approaches use the experimental label information, which is a key component of our method.
-In fact, the only method to our knowledge that leverages such information is the VAE approach developed in Yang et al.
-(2021), which also minimizes a classification loss from the latent space, which we will consider our main baseline.
-For all methods besides SCOT, we use both SNN and OT to perform the final matching in the latent space.
-Note that when combined with OT, our VAE baseline also resembles Cao et al.
-(2022).
-Figure 3: VAE and classifier validation metrics on the CITE-seq dataset.
-Notice that validation cross-entropy inversely tracks the ground truth matching metrics, and thus can be used as a proxy in practical settings where the ground truth is unknown.
-The same pattern does not hold for the VAE (Yang et al., 2021), which we suspect is because reconstruction is largely irrelevant for matching.
-Model Details   For our method and the VAE, we use the general architecture of a linear classification head on top of a suitable encoder.
-The architecture of this encoder is always the same between the VAE and our method.
-For the image dataset, we use the convolutional neural network from Yang et al.
-(2021) as the encoder, and for the CITE-seq dataset, we use fully-connected MLPs for both RNA (top 200 PC’s) and protein data.
-Optimization was performed using Adam with a one-cycle learning rate scheduler.
-We use the standard cross-entropy loss to train our classifiers for propensity score estimation.
-For other methods, we use existing implementations with suggested default settings.
-Matching Details   Both SNN and OT use the Euclidean distance function to determine neighbours and compute the cost matrix, respectively.
-SCOT uses the correlation distance by default, and we found that this resulted in better performance than Euclidean distance.
-We use only a single neighbour for SNN matching, which interestingly resulted in the best performance.
-Both SCOT and OT solve the entropic regularized OT, for these we use a regularization parameter of $0.05$.
-Method MSE Trace
-(Med (Q1, Q3)) (Med (Q1, Q3)) ×$\times 10^{-3}$
-SCOT 0.0354 0.5964
-VAE+SNN 0.0622 3.116
-(0.0571, 0.0676) (2.818, 3.213)
-VAE+OT 0.0324 7.733
-(0.0316, 0.0350) (7.473, 7.794)
-PS+SNN 0.0552 7.924
-(0.0530, 0.0558) (7.569, 9.504)
-PS+OT 0.0316 18.329
-(0.0300, 0.0330) (17.068, 18.987)
-Rand 0.0709 N/A
-(0.0707, 0.0714)
-Table 1: Results on synthetic image data.
-Metrics are computed on a held out test set over 12 groups corresponding to interventions on the latent position, with approximately 1700 observations per group.
-A trace of $10$ for example corresponds to a total weight of $1710\times 10^{-3}\times 1700=$ 17 on the true matching, out of 1700.
-Method FOSCTTM Trace
-(Median (Q1, Q3)) (Median (Q1, Q3))
-SCOT 0.4596 0.0200
-GLUE+SNN 0.4412 0.0362
-GLUE+OT 0.5309 0.0323
-VAE+SNN 0.3816 0.0612
-(0.3760, 0.3822) (0.0588, 0.0634)
-VAE+OT 0.3953 0.0814
-(0.3912, 0.4045) (0.0777, 0.8895)
-PS+SNN 0.3126 0.0941
-(0.3121, 0.3160) (0.0880, 0.0989)
-PS+OT 0.3049 0.1163
-(0.3008, 0.3078) (0.1093, 0.1250)
-Table 2: Results on CITE-seq data.
-Metrics are computed on a held out test set averaged over a total of 45 cell types (groups) with a varying amount of observations per group.
-As such, the average trace can be difficult to interpret (for example, the maximum trace is near $1$, corresponding to a near-perfect matching within a group with only $7$ cells in the test set), but we remark that OT matching on propensity scores outperforms the other methods within each group as well as on average.
-Evaluation Metrics   We report three evaluation metrics.
-The trace metric computes the average mass that the matching matrix places on the true matches (higher is better), and FOSCTTM reports the Fraction Of Samples Closer Than the True Match (FOSCTTM) ((Demetci et al., 2022), (Liu et al., 2019)) (lower is better, 0.5 corresponds to random guessing).
-For synthetic images where we know the ground true latent, we report the MSE to the true latent after matching.
-Full details of these metrics are given in Section D.1.
-Modality Prediction   We also consider one metric that examines whether matched samples are useful for downstream tasks.
-For this, we chose the cross-modality prediction task of predicting protein levels from RNA expression in the CITE-seq dataset.
-To do this, we train a supervised learning model (in our case, a 2-layer MLP with either MSE loss or the unbiased gradients loss in Section 5) using samples generated by different matching matrices $M$.
-As baselines, we also train on pairs from the ground truth matching ($M_{ii}=1$) and random sampling ($M_{ij}=1/n$).
-In view of the performance observed in Table 1 and Table 2, we only considered $M$ resulting from matching with OT using the propensity score and VAE embeddings.
-Method $R^{2}$ (MSE) $R^{2}$ (Unbiased)333 This refers to training the predictor with the unbiased MSE (Equation 11). The evaluation is still using $R^{2}$ as defined by the standard MSE.
-(Med (Q1, Q3)) (Med (Q1, Q3))
-Rand 0.1383 0.1727
-(0.1372, 0.1402) (0.1701, 0.1731)
-VAE+OT 0.1493 0.1142
-(0.1179, 0.1724) (0.0786, 0.1594)
-PS+OT 0.2174 0.2331
-(0.2062, 0.2228) (0.2069, 0.2504)
-True Pairs 0.2243 N/A [^4] The standard and unbiased MSE are equivalent in this case.
-(0.2234, 0.2257)
-Table 3: Modality prediction results on CITE-seq data.
-Entries represent $R^{2}$ computed on a held-out test set consisting of ground-truth pairs, which is computed as $1-\text{MSE}(\hat{x},x)/\text{MSE}(\bar{x},x)$, where $\hat{x}$ are the predicted protein levels, and $\bar{x}$ is the mean value of the validation set.
-Note random sampling, especially with the unbiased gradient loss, achieves a positive $R^{2}$ against $\bar{x}$ because it samples randomly within $t$, which results in learning group-level statistics.
-6.1 Validation Cross-Entropy Monitor
-When the assumptions are satisfied, the success of our propensity score method hinges on how well we can learn the conditional distributions $p(t\mid X_{i}^{(e,t)})$.
-Thus, the validation cross entropy can be used as a tractable metric that can serve as a model selection monitor for the matching performance, which would be unknown in practice.
-The effectiveness of such a monitor is validated in our experiments, even on real-world CITE-seq data, where we see that a lower validation loss typically corresponds to higher matching performance which is not the case for the VAE (Figure 3).
-We suspect that this is precisely due to the additional requirement that the VAE minimize a reconstruction loss, in addition to the classification cross-entropy.
-6.2 Results
-We checkpoint the classifier and VAEs at the lowest validation loss and report their metrics on a held out test set over 10 random seeds.
-Note that this does not necessarily select the embedding model that exhibits optimal matching, but instead the best model to select without validating against ground truth matching metrics, see Figure 3.
-We further used 10 random seeds for training the MLP used for modality prediction.
-scGLUE is trained according to the public implementation, which uses learning rate reduction and early stopping strategies.
-SCOT is a non-iterative approach and thus we directly report its results.
-We found that OT matching on propensity scores consistently outperforms other methods on all metrics, typically followed by SNN matching on propensity scores, or OT matching on VAE embeddings.
-Furthermore, both the VAE and propensity score, which leverage experimental label information, tend to perform better.
-This suggests that using the propensity score as embeddings for matching, and using OT to perform the final matching both independently improve matching performance.
-Curiously, SCOT performs well on the MSE metric for image data, but only places slightly above random in the trace metric.
-This indicates that it matches scenes with similar latent coordinates, without placing any significant weight on the exact ground truth.
-This suggests that exact matches may not be necessary for a method to be useful for downstream tasks.
-Surprisingly, using the propensity score embeddings with OT matching appears to improve generalization in the modality prediction task (Table 3), when using the unbiased MSE over a model trained on ground truth pairings (indeed, we observed that the ground truth model had a lower training loss, but higher test loss).
-This reveals an unexpected benefit of (soft) matching: we can sample from the conditional to minimize the loss Equation 11, which, with a suitable $M$, results in improved generalization compared to the naive MSE.
-This hinges strongly on the quality of $M$—the $M$ resulting from the VAE embeddings results in worse generalization, and in practice the quality of the matching, at least in absolute terms, remains hidden.
-7 Conclusion
-This work presents a simple algorithm for aligning unpaired data from different modalities.
-The method is both very general—only requiring a classifier to be trained on each modality—and highly effective for matching, which we show both theoretically and empirically.
-As a downstream task, we demonstrate the effectiveness of the resulting matchings for cross modality prediction, which leads to better generalization than the ground truth matching on the dataset we evaluated.
-We suspect that this improved generalization is the result of implicitly enforcing invariance to modality specific information, but more work is needed to evaluate the conditions under which this improved generalization occurs.
-Appendix A Cyclic monotonicity
-We can see the monotonicity requirement Equation 8 as the monotonicity of the function with graph $(\pi_{i}^{(1)},\pi_{i}^{(2)})\in[0,1]^{2}$.
-In higher dimensions, we require that the “graph” satisfies the following cyclic monotonicity property (Villani, 2009):
+A.1 Cyclic Monotonicity
+We can see the monotonicity requirement Eq.
+14 as the monotonicity of the function with graph $(\pi_{i}^{(1)},\pi_{i}^{(2)})\in[0,1]^{2}$.
+In higher dimensions, we require that the “graph” satisfies the following cyclic monotonicity property [Villani, 2009]:
 Definition A.1.
@@
 Thus, if the true pairing is uniquely cyclically monotone, we can recover it with OT.
+However, we are unaware of common violations of (A1) that would satisfy cyclic monotonicity.
+Figure 3: OT matching allows for $t$ to have different effects on the modality specific information, here $u_{i}^{(1)}$ and $u_{i}^{(2)}$, as long as they can be written as transformations that preserve the relative order within modalities.
+Exact OT in 1-d always matches according to the relative ordering, and thus exhibits this type of “no crossing” behaviour shown in the figure on the left.
+The figure on the right shows a case where we would fail to correctly match across modalities because of the crossing shown in orange.
 Appendix B Shared Nearest Neighbours Matching
 Using the propensity score distance, we can compute nearest neighbours both within and between the two modalities.
-We follow Cao & Gao (2022) and compute the normalized shared nearest neighbours (SNN) between each pair of observations as the entry of the matching matrix.
+We follow Cao and Gao [2022] and compute the normalized shared nearest neighbours (SNN) between each pair of observations as the entry of the matching matrix.
 For each pair of observations $(\pi_{i}^{(1)},\pi_{j}^{(2)})$, we define four sets:
@@
 ∎
-Appendix C Omitted Proofs
+Appendix C Proofs
 C.1 Proof of Proposition 3.1
 Proof.
-Let $x^{(e,t)}$ denote the observed modality and $z^{(t)},u^{(e)}$ be the unique corresponding latent values. By injectivity,
-$\pi(x^{(e,t)})$ $=P(t|X^{(e,t)}=x^{(e,t)})$
-$=P(t|Z^{(t)}=z^{(t)},U^{(e)}=u^{(e)})$
-$=P(t|Z^{(t)}=z^{(t)})=\pi(z^{(t)}),$ (17)
-for $e=1,2$, since we assumed $U^{(e)}{\perp\!\!\!\perp}t\mid Z^{(t)}$ in Section 3.
-Since this holds pointwise, it shows that $\pi(X^{(1,t})=\pi(X^{(2,t)})=\pi(X^{(t)})=\pi(Z^{(t)})$ as random variables.
-Now, a classical result of Rubin (1974) gives that $Z^{(t)}{\perp\!\!\!\perp}t\mid\pi(Z^{(t)})$, and that for any other function $b$ (a balancing score) such that $Z^{(t)}{\perp\!\!\!\perp}t\mid b(Z^{(t)})$, we have $\pi(Z^{(t)})=g(b(Z^{(t)}))$.
+Let $x^{(e)}$ denote the observed modality and $z,u^{(e)}$ be the unique corresponding latent values. By injectivity,
+$\pi(x^{(e)})$ $=P(t|X^{(e)}=x^{(e)})$
+$=P(t|Z=z,U^{(e)}=u^{(e)})$
+$=P(t|Z=z)=\pi(z),$ (19)
+for $e=1,2$, since we assumed $U^{(e)}{\perp\!\!\!\perp}t\mid Z$.
+Since this holds pointwise, it shows that $\pi(X^{(1})=\pi(X^{(2)})=\pi(Z)$ as random variables.
+Now, a classical result of Rubin [1974] gives that $Z{\perp\!\!\!\perp}t\mid\pi(Z)$, and that for any other function $b$ (a balancing score) such that $Z{\perp\!\!\!\perp}t\mid b(Z)$, we have $\pi(Z)=g(b(Z))$.
 The first property written in information theoretic terms yields,
-$I(t,Z^{(t)}\mid\pi(Z^{(t)}))=I(t,Z^{(t)}\mid\pi(X^{(t)}))=0,$ (18)
-since $\pi(X^{(t)})=\pi(Z^{(t)})$ as random variables, as required. ∎
+$I(t,Z\mid\pi(Z))=I(t,Z\mid\pi(X^{(e)}))=0,$ (20)
+since $\pi(X^{(e)})=\pi(Z^{(t)})$ as random variables, as required. ∎
 C.2 Proof of Proposition 3.2
@@
 Under this convention, the map $h\circ\pi:\mathbb{R}^{d}\to\mathbb{R}^{T}$ is smooth ($\log$ is smooth, and the densities are smooth by assumption).
-Since it is smooth, it cannot be injective if $T<d$ (Spivak, 2018).
+Since it is smooth, it cannot be injective if $T<d$ [Spivak, 2018].
 Finally, since $h$ is bijective, this implies that $\pi$ cannot be injective.
@@
 D.1 Evaluation Metrics
-We can evaluate how well samples are matched using the ground truth provided by our datasets.
+D.1.1 Known Ground Truth
+In the synthetic image and CITE-seq datasets, a ground truth matching is known, and we can evaluate the quality of the synthetic matching directly against the truth.
 In these cases, the dataset sizes are necessarily balanced, so that $n=n_{1}=n_{2}$.
-In each case, the metric is a function of a $n\times n$ matching matrix $M$, which recall we compute within samples with the same $t$.
+In each case, we evaluate the quality of our $n\times n$ matching matrix $M$, which we compute within samples with the same $t$.
 Our reported results are then averaged over each cluster.
-We randomize the order of the datasets before performing the matching to avoid pathological cases.
-Trace Metric   Assuming the given indices correspond to the true matching, we can compute the average weight on correct matches, which is the normalized trace of $M$:
+Note we randomize the order of the datasets before performing the matching to avoid pathologies.
+Trace Metric   Assuming the sample indices correspond to the true matching, we can compute the average weight on correct matches, which is the normalized trace of $M$:
 $\frac{1}{n}\text{Tr}(M)=\frac{1}{n}\sum_{i=1}^{n}M_{ii}.$ (24)
@@
 For example, exactly matching one sample, while adversarially matching dissimiliar samples for the remainder also yields a trace of $1/n$, which is equal to that of a random matching.
-Latent MSE   Because the image dataset is synthetic, we have access to the ground truth latent values that generated the images, $\mathbf{z}=\{z_{i}\}_{i=1}^{n}$.
-We compute matched latents as $M\mathbf{z}$, the mean projection according to the matching matrix.
-Then, to evaluate the quality of the matching, we compute the MSE:
+Latent MSE   On the image dataset, we have access to the ground truth latent values that generated the images, $\mathbf{z}=\{z_{i}\}_{i=1}^{n}$.
+We compute matched latents as $M\mathbf{z}$, the barycentric projection according to the matching matrix.
+Then, to evaluate the quality of the matching in terms of finding similar latents, we compute the MSE:
 $\text{MSE}(M)=\frac{1}{n}\|\mathbf{z}-M\mathbf{z}\|_{2}^{2}.$ (25)
-FOSCTTM   For the CITE-seq dataset, we can use the Fraction Of Samples Closer Than the True Match (FOSCTTM) (Demetci et al., 2022; Liu et al., 2019) as an alternative matching metric.
-First, we distribute the mass of $\mathbf{x}^{(2)}=\{x_{j}\}_{j=1}^{n}$ to $\mathbf{x}^{(1)}=\{x_{i}\}_{i=1}^{n}$ as $\hat{\mathbf{x}}^{(1)}=M\mathbf{x}^{(2)}$, resulting in a projection of the first modality $\mathbf{x}^{(1)}$ in the space of the second modality.
+FOSCTTM   We do not have access to ground truth latents in the CITE-seq dataset, so use the Fraction Of Samples Closer Than the True Match (FOSCTTM) [Demetci et al., 2022, Liu et al., 2019] as an alternative matching metric.
+First, we use $M$ to project $\mathbf{x}^{(2)}=\{x_{j}\}_{j=1}^{n}$ to $\mathbf{x}^{(1)}=\{x_{i}\}_{i=1}^{n}$ as $\hat{\mathbf{x}}^{(1)}=M\mathbf{x}^{(2)}$.
 Then, we can compute a cross-modality distance as follows.
@@
 $+\sum_{j=1}^{n}\bigg{(}\frac{1}{n}\sum_{i\neq j}\mathds{1}\{d(\mathbf{x}_{j}^{(1)},\hat{\mathbf{x}}_{i}^{(1)})<d(\mathbf{x}^{(1)}_{j},\hat{\mathbf{x}}_{j}^{(1)})\}\bigg{)}\bigg{]},$ (27)
-where notice that this is a function of $M$ through the computation $\hat{\mathbf{x}}^{(1)}=M\mathbf{x}^{(2)}$.
+where notice that this evaluates $M$ through the computation $\hat{\mathbf{x}}^{(1)}=M\mathbf{x}^{(2)}$.
 As a baseline, we should expect a random matching, when distances between points are randomly distributed, to have an FOSCTTM of $0.5$.
+Prediction Accuracy   We also trained a cross-modality prediction (translation) model $f_{\theta,M}$ to predict CITE-seq protein levels from gene expression based on matched pseudosamples.
+Let $\mathbf{x^{(1)}}=\{x_{i}\}$, $\mathbf{x^{(2)}}=\{x_{j}\}$ denote protein and gene expression, respectively.
+We trained a simple 2-layer MLP minimizing either the standard MSE, using pairs $(x_{i},\hat{x}_{j})$, $\hat{x}_{j}\sim M_{i\cdot}$, or following the projected loss with unbiased estimates in Section 5.
+Each batch in general consists of samples from all $t$, but the $\hat{x}_{j}$ sampling step occurs within the perturbation.
+Let $\hat{\mathbf{x}}^{(1)}_{test}=\{f_{\theta,M}(x_{j})\}$.
+We report the $R^{2}$ on a randomly held-out test set of ground truth pairs (again, consisting of samples from all $t$), which is defined as the following:
+$R^{2}(f_{\theta,M})=\frac{MSE(\mathbf{x}^{(1)}_{test},\hat{\mathbf{x}}^{(1)}_{test})}{MSE(\mathbf{x}^{(1)}_{test},\bar{\mathbf{x}}^{(1)}_{test})},$ (28)
+where $\bar{\mathbf{x}}^{(1)}_{test}$ is the naive mean (over all perturbations) estimator which acts as a baseline.
+D.1.2 Unknown Ground Truth
+We train a cross-modality prediction model to predict gene expression from cell images based on matched pseudosamples in the same way as in CITE-seq, but only using the projected loss with unbiased estimates.
+Denote this model for a matching matrix $M$ by $f_{\theta,M}$.
+Because we do not have access to ground truth pairs within each perturbation, we resort to distributional metrics.
+Let $\mathbf{x^{(1)}}_{t}=\{x_{i,t}\}_{i=1}^{n_{t1}}$, $\mathbf{x^{(2)}}_{t}=\{x_{j,t}\}_{j=1}^{n_{t2}}$ denote gene expression and cell images in a held out test set respectively in perturbation $t$.
+Let $\hat{\mathbf{x}}^{(1)}_{t}=\{f_{\theta,M}(x_{j,t})\}_{j=1}^{n_{t2}}$.
+We compute empirical versions of statistical divergences
+$D_{t}(f_{\theta,M}):=D(\mathbf{x}^{(1)}_{t},\hat{\mathbf{x}}^{(1)}_{t}),$ (29)
+where $D$ is either the L2 norm of the difference in empirical mean, empirical Kullback-Leibler divergence or 1-Wasserstein distance.
+We report these weighted averages of $D_{t}$ over the perturbations $t$ according to the number of samples in the modality of prediction interest.
 D.2 Models
-In this section we describe experimental details pertaining to the propensity score and VAE (Yang et al., 2021).
-SCOT (Demetci et al., 2022) and scGLUE (Cao & Gao, 2022) are used according to tutorials and recommended default settings by the authors.
-Loss Functions   The propensity score approach minimizes the standard cross-entropy loss for both modalities, as implemented in PyTorch 2.0.1.
+In this section we describe experimental details pertaining to the propensity score and VAE [Yang et al., 2021].
+SCOT [Demetci et al., 2022] and scGLUE [Cao and Gao, 2022] are used according to tutorials and recommended default settings by the authors.
+Loss Functions   The propensity score approach minimizes the standard cross-entropy loss for both modalities.
 The VAE includes, in addition to the standard ELBO loss (with parameter $\lambda$ on the KL term), two cross-entropy losses based on classifiers from the latent space: one, weighted by a parameter $\alpha$ to classify $t$ as in the propensity score, and another, weighted by a parameter $\beta$, that classifies which modality the latent point belongs to.
@@
 We follow Yang et al.
-(2021) and set $\alpha=1$, $\beta=0.1$, but found that $\lambda=10^{-9}$ (compared to $\lambda=10^{-7}$ in Yang et al.
-(2021)) resulted in better performance.
+[2021] and set $\alpha=1$, $\beta=0.1$, but found that $\lambda=10^{-9}$ (compared to $\lambda=10^{-7}$ in Yang et al.
+[2021]) resulted in better performance.
 We used batch size 256 in both instances and trained for either 100 epochs (image) or 250 epochs (CITE-seq).
+For the VAE and classifiers of experiment 3, we use an Adam optimizer with learning 0.001 and weight decay 0.001 and max epoch of 100 (PerturbSeq) and 250 (single cell images) using batch sizes of 256 and 2048 correspondingly.
+We follow similar settings as Yang et al.
+[2021] and implement $\alpha=1$ with $\lambda=10^{-9}$, and since we do not have matched data, $\beta=0$.
+For the cross-modal prediction models in experiment 3, we use Stochastic Gradient Descent optimizer with learning rate 0.001 and weight decay 0.001 with max epochs 250 and batch size 256.
+We implement early stopping with delay of 50 epochs which we then checkpoint the last model to use for downstream tasks
 Architecture   For the synthetic image dataset, we use an 5-layer convolutional network (channels $=32,54,128,256,512$) with batch normalization and leaky ReLU activations, with linear heads for classification (propensity score and VAE) and posterior mean and variance estimation (VAE).
@@
 For the CITE-seq dataset, we use a 5-layer MLP with constant hidden dimension $1024$, with batch normalization and ReLU activations (adapted from the fully connected VAE in Yang et al.
-(2021)) as both the encoder and VAE decoder.
+[2021]) as both the encoder and VAE decoder.
 We use the same architecture for both modalities, RNA-seq (as we process the top 200 PCs) and protein.
-Optimal Transport   We used POT (Flamary et al., 2021) to solve the entropic OT problem, using the log-sinkhorn solver, with regularization strength $\gamma=0.05$.
+For the PerturbSeq classifier encoder, we use a 2-layer MLP architecture.
+Each layer consists of a linear layer with an output feature dimension of 64, followed by Rectified Linear Unit (ReLU) activation, Batch Normalization, and dropout (p=0.1).
+A final layer with Leaky ReLU activation that brings dimensionality to 128 before feeding into a linear classification head with an output feature dimension of 25.
+For the single-cell image encoder classifier, we use a proprietary Masked Autoencoder [Kraus et al., 2023] to generate 1024-dimensional embeddings.
+Subsequently, a 2-layer MLP is trained on these embeddings.
+Each MLP layer has a linear layer, Batch Normalization, and Leaky ReLU activation.
+The output feature dimensions of the linear layers are 512 and 256, respectively, and the latent dimension remains at 1024 before entering a linear classification head with an output feature dimension of 25.
+Optimal Transport   We used POT [Flamary et al., 2021] to solve the entropic OT problem, using the log-sinkhorn solver, with regularization strength $\gamma=0.05$.
 D.3 Data
-Synthetic Data   We follow the data generating process Section 3 to generate coloured scenes of two simple objects (circles, or squares) in various orientations and with various backgrounds.
+Synthetic Data   We follow the data generating process Eq.
+1 to generate coloured scenes of two simple objects (circles, or squares) in various orientations and with various backgrounds.
 The position of the objects are encoded in the latent variable $z$, which is perturbed by a do-intervention (setting to a fixed value) randomly sampled for each $t$.
@@
 Scenes are generated using a rendering engine from PyGame as $f^{(e)}$.
-Example images are given in Figure 4.
+Example images are given in Fig.
+4.
 Figure 4: Example pair of synthetic images with the same underlying $z$.
-CITE-seq Data
-We also use the CITE-seq dataset from Lance et al.
-(2022) as a real-world benchmark (obtained from GEO accession GSE194122).
+CITE-seq Data   We also use the CITE-seq dataset from Lance et al.
+[2022] as a real-world benchmark (obtained from GEO accession GSE194122).
 These consist of paired RNA-seq and surface level protein measurements, and their cell type annotations over $45$ different cell types.
@@
 For more details, see Lance et al.
-(2022).
+[2022].
+PerturbSeq and Single Cell Image Data   We collect single-cell PerturbSeq data (200 genes) and single-cell painting images in HUVEC cells with 24 gene perturbations and a control perturbation, resulting in 25 labels for matching across both modalities.
+The target gene perturbations are selected based on the 24 genes with the highest number of cells affected by the CRISPR guide RNAs targeting those genes.
+The PerturbSeq data is filtered to include the top 200 genes with the highest mean count, then normalized and log-transformed.
+The single-cell painting images are derived from multi-cell images, with each single-cell nucleus centered within a 32x32 pixel box.
+We use scVI Lopez et al.
+[2018] to embed the raw PerturbSeq counts into a 128-dimensional space before training the gene expression classifier.
+Similarly, we train our image classifier using 1024-dimensional embeddings obtained from a pre-trained Masked Autoencoder Kraus et al.
+[2023], He et al.
+[2022].
+Following matching, we perform cross-modality translation from the single-cell embeddings to the transformed gene expression counts.
+D.4 Supplementary Results
+In Distribution Out of Distribution
+Method Wasserstein-1 ($\downarrow$) L2 Norm ($\downarrow$) Wasserstein-1 ($\downarrow$) L2 Norm ($\downarrow$)
+Med (Q1, Q3) Med (Q1, Q3)
+PS+OT 4.199 3.280 5.394 7.219
+(4.173, 4.226) (3.267, 3.284)
+VAE+OT 4.339 3.490 5.629 7.444
+(4.314, 4.348) (3.486, 3.495)
+Random 4.499 3.826 6.239 7.793
+(4.478, 4.525) (3.823, 3.828)
+Table 3: Wasserstein-1 and L2 norm distance values for PerturbSeq and single cell images experiments where distance is evaluated between cross-modal predictions and actual gene expression values.